### v1

# DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD

Abstract
Distributed machine learning in high end-to-end latency and low, varying bandwidth network environments undergoes severe throughput degradation. Due to its low communication requirements, distributed SGD (D-SGD) remains the mainstream optimizer in such challenging networks, but it still suffers from significant throughput reduction. To mitigate these limitations, existing approaches typically employ gradient compression and delayed aggregation to alleviate low bandwidth and high latency, respectively. To address both challenges simultaneously, these strategies are often combined, introducing a complex three-way trade-off among compression ratio, staleness (delayed synchronization steps), and model convergence rate. To achieve the balance under varying bandwidth conditions, an adaptive policy is required to dynamically adjust these parameters. Unfortunately, existing works rely on static heuristic strategies due to the lack of theoretical guidance, which prevents them from achieving this goal. This study fills in this theoretical gap by introducing a new theoretical tool, decomposing the joint optimization problem into a traditional convergence rate analysis with multiple analyzable noise terms. We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression on training performance, filling a critical gap in understanding how compressed and delayed gradients affect training. Furthermore, by integrating the convergence rate with a network-aware time minimization condition, we propose DeCo-SGD, which dynamically adjusts the compression ratio and staleness based on the real-time network condition and training task. DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and static strategy in high-latency and low, varying bandwidth networks, respectively.
1 Introduction
Recent breakthroughs in large language models (LLMs) (e.g., GPT-4 gpt-4 , DeepSeek-v3 deepseek ) have been accompanied by an explosive growth in model parameters, leading to unprecedented communication demands for distributed training scalalaws2 ; ScalaLaws1 . State-of-the-art (SOTA) distributed training strategies flashattention ; megatron-lm2 are predominantly optimized for high-speed interconnects like PCIe 5.0 and NVLink nvidia , where bandwidth exceeds $10$ Gbps and latency remains under $10$ ms. Although these assumptions are reasonable in well-resourced data centers, they often fail to hold in dynamic network environments with high latency and low bandwidth jaghouar2024opendiloco . Unfortunately, such unfavorable network conditions are common in distributed training over wide area networks (WAN), such as federated learning dagctmc and cross-geographic distributed learning hsieh2017gaia . WAN-based distributed training lowers the barrier for LLM pre-training, empowering individual researchers and small labs while reducing the reliance on costly, specialized networking SIGCOMM24StellaTrain .
In these challenging dynamic network settings, distributed SGD (D-SGD) and its variants, such as FedAVG fedavg and D-QSGD cui2021slashing , remain the dominant optimization methods due to their communication-friendly nature. Despite its low communication requirements, D-SGD still suffers from significant throughput degradation under high-latency and low-bandwidth conditions. As illustrated in Fig. 1, when the bandwidth drops below $2$ Gbps and the latency exceeds $200$ ms, the throughput of D-SGD in training GPT-2 with four nodes reduces to $50\%$ of maximum achievable throughput [^1] The maximum achievable throughput is determined by the compute capacity of the machines, without network constraints.. {wrapfigure}t0.4 [Uncaptioned image]
[…]
Gradient compression and delayed aggregation are common strategies to improve the throughput of D-SGD in a lossy manner, but joint usage introduces a complex three-way trade-off. Gradient compression [^2] In this work, the gradient compression algorithm carries vanilla error-feedback (EF) stich2019error by default. EF is a mechanism that collects and reuses the errors from the gradient compression to retain the model accuracy. Other EF strategies, such as EF21 EF21 or EControl Stich2023EControl are out of this paper’s scope. reduces communication overhead by compressing gradient information before transmission, to mitigate the low bandwidth problem dgc . Delayed aggregation DGA applies updates using gradients computed several iterations earlier to alleviate the high latency problem. To overcome the problem of low-bandwidth and high-latency at the same time in WANs, we need to use these two strategies jointly, introducing intricate trade-offs in two key dimensions: (1) Iteration-to-accuracy: The impact of compressing delayed gradients on convergence remains uncertain, as it is unclear to what extent aggregating delayed compressed gradients will affect model accuracy, due to the lack of theoretical analysis in prior work. (2) Time-to-iteration: The effect of staleness and compression ratio on communication-computation overlap is not fully understood. Specifically, the existence of a threshold compression ratio, below which further compression no longer improves per-iteration time, remains unexplored. This threshold varies dynamically with the bandwidth and staleness.
To achieve optimal time-to-accuracy, we analyze the above two dimensions separately. First, in iteration-to-accuracy, we propose a new analytical tool called the Nested Virtual Sequence. This method builds on the original virtual sequence by constructing multiple virtual queues, breaking down the complex noise term into multiple analyzable noise terms. With this tool, we are the first to derive the convergence rate of distributed SGD with delayed aggregation and gradient compression (DD-EF-SGD). This convergence rate establishes the iteration-to-accuracy relationship and reveals that the value of staleness exponentially worsens the negative effect of the gradient compression on the model convergence. Second, we use mathematical induction to model the relationship between the end-to-end time and the number of iterations in DD-EF-SGD. In this way, we build the time-to-accuracy relationship and propose an adaptive algorithm (shown in Fig. 3) that can select the optimal compression ratio and staleness in the challenging dynamic network conditions, thus achieving the balanced three-way trade-off during the whole training.
In a word, the contributions of our work are summarized as below:
  • We propose the Nested Virtual Sequence, a theoretical tool to analyze complex distributed SGD variants, by reducing it to a standard D-SGD process and several analytical noise terms. We are the first to derive the convergence rate of DD-EF-SGD with the tool, revealing that the delayed aggregation exponentially worsens the negative effect of the compression on the training process.
  • We mathematically establish an approximate relationship between end-to-end time and accuracy. Based on this, we propose DeCo-SGD, a joint optimization algorithm that adaptively calculates the optimal compression ratio and staleness according to the training environments.
  • We apply DeCo-SGD  in challenging networks environments (average bandwidth < $1$Gbps, latency $\geq$ $100$ms) across diverse model architectures including CNN, GPT and ViT. Experimental results validate the great three-way trade-offs of our design, achieving up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and the SOTA static strategy, respectively.
2 Preliminaries
We specifically concentrate on distributed machine learning (DML) in data-parallel mode using the DD-EF-SGD as the optimizer.
Figure 2: The running timelines for D-SGD and D-SGD variants. Both D-SGD and D-EF-SGD are serial processes, with D-EF-SGD achieving a reduced transmission time. Similarly, DD-SGD and DD-EF-SGD operate serially; while DD-SGD maintains the same communication time as D-SGD, DD-EF-SGD reduces transmission time.
2.1 The Optimization Problem of DML
In the data-parallel paradigm, the training dataset is partitioned across $n$ worker nodes, each holding a subset of the data. The optimization problem of DML aims to minimize a global loss function $f(\cdot)$ based on the model parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $f(\cdot)$ is defined as a weighted sum of local loss functions of the $i$-th worker $f_{i}(\cdot)$:
$\underset{\mathbf{x}}{\min}\left[f(\mathbf{x}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})\right].$
2.2 D-SGD and Its Variants
2.2.1 D-SGD
At $t$-th iteration, the $i$-th worker computes gradients $\mathbf{g}_{t}^{i}$ based on its local data and model parameters at $t$-th iteration $\mathbf{x}_{t}$. $\gamma$ is the stepsize and these local gradients are aggregated to update the global model, following $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. As shown in Fig. 2, D-SGD is a serial computation process. In this process, computation and communication take place sequentially in each iteration, and communication includes end-to-end latency and transmission time.
2.2.2 D-SGD with Gradient Compression and EF (D-EF-SGD)
D-EF-SGD stich2020communication enhances D-SGD with the sparsification gradient compression [^3] In this work, the sparsification compressor in D-EF-SGD is Top-$k$ by default, a well-performed compressor in many hybrid gradient compressors stc ; CocktailSGD . by incorporating EF mechanisms to compensate for sparsification compression errors. Sparsification compression is $\mathbf{e}_{t}^{i}$ represents the error term of the $i$-th worker at the $t$-th iteration, storing the discrepancy between the compressed gradient and the true gradient. We denote $\mathbf{C}_{\delta}(\cdot)$ as the compression operator with the compression ratio $\delta$, $\Delta_{t}^{i}$ as the compressed update at $t$-th iteration in $i$-th worker, and the update rules for D-EF-SGD are as follows:
$\begin{gathered}\Delta_{t}^{i}=\mathbf{C}_{\delta}(\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}),\quad\quad\quad\mathbf{e}_{t+1}^{i}=\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}-\Delta_{t}^{i},\quad\quad\quad\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\Delta_{t}^{i}.\end{gathered}$
As shown in Fig. 2, D-EF-SGD retains a serial process, but reduces transmission time compared to D-SGD.
2.2.3 D-SGD with Delay Aggregation (DD-SGD)
DD-SGD DGA addresses communication delays by accounting for the delay staleness $\tau$ between the computation of gradients and their application to the model update stich2019error . $\mathbf{g}_{t-\tau}^{i}$ represents the gradient computed by worker $i$ at the $(t-\tau)$-th iteration and the update rule shows as $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}.$ The delay aggregation does not change the communication time for each iteration. As shown in Fig. 2, DD-SGD parallelizes computation and communication by performing the $t$-th computation at the $t$-th iteration and sending $t$-th gradient, but the aggregation applies the gradient computed at $(t-\tau)$-th iteration. When combined with gradient compression (DD-EF-SGD), the transmission time is further reduced, as shown in the last part of Fig. 2.
2.3 Virtual Sequence
Virtual Sequence stich2018sparsified ; mania2017perturbed is an analytical tool that simplifies the convergence analysis of D-SGD variants like D-EF-SGD. It defines an auxiliary sequence, $\tilde{\mathbf{x}}_{t}$, which evolves according to an idealized update rule and serves as a reference for the actual iterates, $\mathbf{x}_{t}$, accounting for noise introduced by compression and delay. We define the noise term as $B_{t}$, then we have $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$. We generalize updates as $\mathbf{v}_{t}$, (i.e., $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\mathbf{v}_{t}$) and $B_{t}$ can be computed in $B_{t+1}=B_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\mathbf{v}_{t}$. Then we have $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\textbf{g}_{t}^{i}$. This can help theoretically analyze D-SGD variants, but fails in complex variants like DD-EF-SGD, because the noise term decoupled from DD-EF-SGD cannot be analyzed. A new method is needed to analyze these complex D-SGD variants.
3 Theoretical Analysis of DD-EF-SGD
We derive the convergence rate of DD-EF-SGD. Notation list and detailed proof are in the Appendix.
3.1 Regular Assumptions
We assume functions are $L$-smooth with the gradient noise of D-SGD presumed to exhibit zero mean and a variance of $\sigma^{2}$. We gauge data heterogeneity using constants $\zeta_{i}^{2}>0$, which cap the variance among the $n$ workers. In Theorem 2, we posit that objective functions are $\mu$-strongly convex. Detailed assumptions are shown below.
Assumption 1 ($L$-smoothness). We assume $L$-smoothness of $f_{i},\ i\in[n]$, that is for all $\mathbf{x},\ \textbf{y}\in\mathbb{R}^{d}$:
$\lVert\nabla f_{i}(\textbf{y})-\nabla f_{i}(\mathbf{x})\rVert\leq L\lVert\textbf{y}-\mathbf{x}\rVert.$ (1)
Assumption 2 ($\mu$-strongly convexity). We assume $\mu$-strong convexity of $f_{i},i\in[n]$, that is for all $\mathbf{x},\textbf{y}\in\mathbb{R}^{d}$:
$f_{i}(\mathbf{x})-f_{i}(\textbf{y})\geq\langle\nabla f_{i}(\textbf{y}),\mathbf{x}-\textbf{y}\rangle+\frac{\mu}{2}\lVert\nabla f_{i}(\mathbf{x})-\nabla f_{i}(\textbf{y})\rVert^{2}.$ (2)
Assumption 3 (Bounded gradient noise). We assume that we have access to stochastic gradient oracles $\mathbf{g}^{i}_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for each $f_{i},\ i\in[n]$. We define the stochastic gradient noise of $i$-th worker as $\xi^{i}$ and for simplicity, we only consider the instructive case of uniformly bounded $\xi^{i}$ for all $\mathbf{x}\in\mathbb{R}^{d},\ i\in[n]$:
$\mathbf{g}^{i}_{t}=\nabla f_{i}(\mathbf{x}_{t})+\boldsymbol{\xi}^{i},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\boldsymbol{\xi}^{i}=\textbf{0}_{d},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\lVert\boldsymbol{\xi}^{i}\rVert^{2}\leq\sigma^{2}.$ (3)
Assumption 4 (Measurement of data heterogeneity). We measure data dissimilarity by constants $\zeta_{i}^{2}\geq 0$ that bound the variance across the $n$ nodes. We have:
$\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+\lVert\nabla f(\mathbf{x})\rVert^{2},\quad\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2},\quad\forall\mathbf{x}\in\mathbb{R}^{d},i\in[n].$ (4)
These assumptions are standard, following previous works cui2022infocom ; stich2020communication . For Assumption 3, we can assume that the $\sigma^{2}$ increases with the gradient norm hardthreshold . In this way, by replacing $\sigma$ with a new $\sigma^{\prime}$, one can obtain a tighter bound. Similarly in Assumption 4, the general version is $\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+Z\lVert\nabla f(\mathbf{x})\rVert^{2}$, and we take $Z=1$ by default. One can take a more general bound by adding $Z$. However, extending the analysis to cover these generalized assumptions is beyond the scope of this work.
3.2 Nested Virtual Sequences (NVS)
To solve the problem that the traditional Virtual Sequence method cannot analyze complex D-SGD variants like DD-EF-SGD. We propose NVS, which is composed of two parts. In the first part, we have
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\mathbf{v}_{t},\quad\quad\quad B_{t+1}=B_{t}+\tilde{\mathbf{v}}_{t}-\mathbf{v}_{t}.$ (5)
Compared with Virtual Sequence, we change the second term in Eq. 5 from the update $\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$ into $\tilde{\mathbf{v}}_{t}$. We define the first virtual sequence $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-B_{t}$ and observe that $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\mathbf{v}_{t}-B_{t}-\tilde{\mathbf{v}}_{t}+\mathbf{v}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t}$. In the second part, we use the virtual sequence framework based on the new items $\tilde{\mathbf{x}}_{t}$ and $\tilde{\mathbf{v}}_{t}$. We define the second noise term as $\tilde{B}_{t}$ and have:
$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\tilde{B}_{t+1}=\tilde{B}_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\hat{\mathbf{x}}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{B}_{t}.$
Similarly, it is easy to get $\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. In this way, we decouple DD-EF-SGD into a standard D-SGD process and two analyzable noise terms:
$\mathbf{v}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{C}_{\delta}(\mathbf{g}_{t-\tau}^{i}+e_{t-\tau}^{i}),\quad B_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}e_{t-\tau}^{i},$
$\tilde{\mathbf{v}}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i},\quad\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-j}^{i}.$ (6)
3.3 Convergence Rate of DD-EF-SGD
Theorem 1 (Non-convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{1}{\epsilon}+\frac{\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$ (7)
iterations of DD-EF-SGD, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total number of iterations, chosen at random uniformly.
Remark 1.
($\phi$ determining the convergence in the non-degradation condition) In typical DML scenarios, $\frac{\tau}{n}\leq 1$. As the training scale increases, $\tau$ tends to decrease due to the larger model requiring larger computation time, and $n$ tends to increase, resulting in an even smaller $\frac{\tau}{n}$. While to reduce the communication traffic, $\delta$ tends to be $10\%$ or less, common in DML CocktailSGD ; grace , leading to the order of magnitude of $\phi$ larger than that of $\frac{\tau}{n}$. Then the convergence rate is $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}+\frac{\sqrt{\frac{\phi}{\delta}}+\sqrt{\tau}}{\epsilon})$. Similar to prior works dagctmc , the frist term $\frac{\sigma^{2}}{n\epsilon^{2}}$ is independent of $\delta$ and $\tau$. The second term $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$ has a greater impact on the overall convergence rate compared to the third term.
Let us focus on $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$, in LLMs pre-training scenarios, local datasets are typically allocated centrally, resulting in low data heterogeneity $\zeta$. Moreover, due to the small batch sizes commonly used in LLM training (limited by the constraint of GPU memory), $\sigma$ tends to be large. Under such circumstances, $\frac{\zeta^{2}}{\delta}\ll\sigma^{2}$ and $\phi$ captures the convergence rate. While in Federated Learning (large $\zeta$) or training small models (small $\sigma^{2}$), the key factor is $\phi^{\prime}=\frac{1-\delta}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}$. In this work, we focus on the first scenario where $\phi$ is the key factor that determines the convergence of DD-EF-SGD. For the analysis in Federated Learning or training small models, simply replacing $\phi$ with $\phi^{\prime}$ is enough and we do not discuss this later in the paper.
Remark 2.
(Degradation condition) DD-EF-SGD has two degradation conditions, namely, without compression ($\delta=1$) and without delayed aggregation ($\tau=0$). When $\delta=1$, $\phi=0$ and DD-EF-SGD degrades into DD-SGD. Then the convergence rate is equal to $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\sqrt{\tau}}{\epsilon})$, same with the convergence rate of DD-SGD stich2022sharper ; stich2019error . When $\tau=0$, DD-EF-SGD degrades to D-EF-SGD, and $\phi=\frac{1-\delta}{\delta}$. Then the convergence rate becomes $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\zeta/\sqrt{\delta}+\sigma}{\sqrt{\delta}\epsilon^{3/2}}+\frac{L}{\delta\epsilon})$, same with the convergence rate of D-EF-SGD stich2020communication ; dagc .
Theorem 2 (Convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$ (8)
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
Remark 3.
Same to remarks in Theorem 1, in the degradation conditions, the convex convergence rate is the same as that of DD-SGD (under $\delta=1$) and D-EF-SGD (under $\tau=0$). In the non-degradation conditions, the convergence rate can be written into $\mathcal{O}(\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{\sqrt{\frac{L\phi}{\delta}}+\sqrt{L\tau}}{\mu})$, $\phi$ dominates the convergence rate of DD-EF-SGD in the convex case, too.
4 Mathematical Modeling Time to Accuracy in DD-EF-SGD
[…]
4.1 Time to Iteration in DD-EF-SGD
Theorem 3 (Estimation of $T_{avg}$ in DD-EF-SGD).
Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed. Then the average iteration time $T_{avg}$ satisfies the following approximation:
$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\delta\cdot S_{g}/a}{\tau+1},\frac{\delta\cdot S_{g}}{a},T_{\text{comp}}\right\}},$ (9)
where the approximation error is upper bounded by $O(\frac{1}{t})$.
Remark 4 (The optimal compression ratio $\delta^{*}$).
We can minimize $T_{avg}$ to $T_{comp}$ by enlarging $\tau$ and choosing a small $\delta$ Let $\delta^{*}(\tau)=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$, we can obtain that $\delta\leq\delta^{*}(\tau)$. Given $\tau\geq\frac{b}{T_{comp}}$, $\delta^{*}=\delta^{*}(\tau)$ This is due to the fact that the more aggressive $\delta$ cannot get the performance gain, which is covered by the delay while leading to accuracy degradation. In this way, $\delta^{*}$ is the maximal value in the given range.
4.2 DeCo: Joint Optimization for Delay Staleness and Gradient Compression in DML
Figure 3: The high level design of DeCo-SGD. Our design aims to minimize the end-to-end time while achieving the target accuracy. DeCo-SGD adaptively adjusts the compression ratio $\delta$ and delay staleness $\tau$ based on the dynamic network conditions monitored by the worker.
Input: $S_{g}$, $a$, $b$, $T_{comp}$
Output: Optimal $\tau$, $\delta$ with minimal $\phi$
$\phi_{min}=\infty$
for $\tau\in\{\lceil\frac{b+S_{g}/a}{T_{comp}}\rceil,\cdots,1+\lceil\frac{b}{T_{comp}}\rceil,\lceil\frac{b}{T_{comp}}\rceil\}$  do
      $\delta=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$;
     /* Calculate the value of $\phi$/
      $\phi=\frac{1-\delta}{\delta\cdot(1-\frac{\delta}{2})^{\tau}}$;
     if $\phi\leq\phi_{min}$ then
           $\phi_{min}=\phi$;
           $\tau^{*},\delta^{*}=\tau,\delta$;
           end if
           end for
          Return $\tau^{*}$, $\delta^{*}$;
Algorithm 1 DeCo
Input: $n$, $\gamma$, $S_{g}$, $T_{comp}$, initial parameters $\textbf{x}_{0}$, initial local error $\textbf{e}_{0}^{i}=\textbf{0}_{d}$, the frequency to update DeCo $E$
Output: $\textbf{x}_{T}$
for $i=1,\ldots,n$ do
       if $t\mod E==1$ then
             /* Update $\tau$ and $\delta$ with DeCo per $E$ iterations */
             Get $a$, $b$ from the network;
             $\tau$, $\delta$ = DeCo($S_{g}$, $a$, $b$, $T_{comp}$);
       end if
      $\hat{\Delta}_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i})$;
       $\textbf{e}_{t+1-\tau}^{i}=\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i}-\hat{\Delta}_{t}^{i}$;
       $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\hat{\Delta}_{t}^{i}$;
end for
Return $\textbf{x}_{T}$;
Algorithm 2 DeCo-SGD
Our design goal is to get the minimal end-to-end time to achieve a specific accuracy, which is equivalent to solving the minimum of the binary function $\phi$. Since $\tau$ is discrete, the minimum of $\phi$ cannot be directly solved by the partial derivatives. We use the critical condition $T_{avg}=T_{comp}$ (i.e., there are no redundant bubbles in the pipeline) to reduce the global optima into a countable number of local optima. The mathematical description of this goal is
$\min\ \phi(\delta,\tau)=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}},\quad\textit{s.t.}\ T_{avg}=T_{comp},\quad\tau\in\mathbb{N},\quad\delta\in\left(0,1\right].$ (10)
Combined with Remark 4, we can rewrite Eq. 10 into:
$\min\ \phi(\delta^{*}(\tau),\tau)=\frac{1-\delta^{*}(\tau)}{\delta^{*}(\tau)(1-\frac{\delta^{*}(\tau)}{2})^{\tau}}\quad\textit{s.t.}\ \tau\in\left[\lceil\frac{b}{T_{comp}}\rceil,\cdots,\lceil\frac{b+\frac{S_{g}}{a}}{T_{comp}}\rceil\right].$ (11)
We traverse the range of $\tau$ to get $\phi^{*}$ to achieve the fastest convergence rate and choose the smallest $\tau$ as $\tau^{*}$ The procedure for obtaining $\tau^{*}$ and $\delta^{*}$ is referred to as the DeCo algorithm (shown in Algo. 1).
The implementation of DeCo in D-SGD is called DeCo-SGD (pseudo-code shown in Algo. 2 and design shown in Fig. 3). DeCo-SGD sets a frequency $E$ to update DeCo. $E$ describes the sensitivity of DeCo. To enable frequent adjustments of the compression ratio and delay staleness in response to a dynamic network environment, a smaller $E$ value can be adopted, such as $E=1$. In the normal case, a middle $E$ like $100$ is enough. DeCo-SGD incurs a negligible extra cost compared to vanilla DD-EF-SGD and the extra time complexity is $\mathcal{O}(\frac{T}{E})$. The increase in $n$ does not result in additional cost.
5 Evaluation Experiments
5.1 Experimental Settings
We evaluate our approach on diverse CV and NLP tasks using CNN lecun1998gradient , ViT vit , and GPT gpt2 models, including FashionMNIST xiao2017fashion , CIFAR-$10$, ImageNet imagenet , and Wikitext wikitext . Our baselines include D-SGD, Accordion accordion , DGA DGA , and CocktailSGD CocktailSGD . CocktailSGD is a SOTA static strategy that combines gradient compression and delayed aggregation using fixed parameters. If not specified, all compressors are Top-$k$ gradient compression dagctmc , except for CocktailSGD.
5.2 Comparison of Training Speed
As illustrated in Fig. 4, DeCo-SGD demonstrates significant improvements in training speed compared to D-SGD, Accordion, DGA, and CocktailSGD across all tasks. All experiments were conducted with $4$ worker nodes. Take ViT@ImageNet as an example, when the accuracy reaches 70%, DeCo-SGD achieves a 4.99×4.$9$ speedup444speed-up is defined as the ratio of the baseline method’s training time to that of DeCo-SGD. over D-SGD and a 1.34×1.34$\times$1.34 × speedup over CocktailSGD. At 80% accuracy, the speedup increases to 5.01×5.01$\times$5.01 × and 1.37×1.37$\times$1.37 ×, respectively. Compared with other baseline methods, the benefits of DeCo-SGD are particularly obvious, mainly for two reasons: (1) their parameter settings are suboptimal; (2) they cannot adapt to the dynamic network environment in real time. In contrast, DeCo-SGD utilizes theoretical guidance and runtime feedback to jointly optimize compression and latency, thereby achieving a more effective communication-computing trade-off throughout the training process.
Figure 4: Training time comparison across different model-dataset pairs.
5.3 Scalability Verification of DeCo-SGD
DeCo-SGD is inherently scalable, as the computational complexity of DeCo-SGD in finding $\delta^{*}$ and $\tau^{*}$ is theoretically independent of the number of worker nodes. Fig. 5 illustrates the performance of DeCo-SGD across different numbers of worker nodes, scaling from $n=4$ to $n=32$. All experiments were conducted under a fixed latency of $200$ms, with bandwidth fluctuating around $100$Mbps.
The results demonstrate that DeCo-SGD consistently maintains performance improvements as the number of nodes increases. Specifically, for GPT@Wikitext, DeCo-SGD achieves up to 3.77×3.$7$ and 1.21×1.21$\times$1.21 × speed-up over D-SGD and CocktailSGD, respectively, even at $n=32$. Similarly, for ViT@ImageNet, DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over CocktailSGD. These results highlight the adaptability of DeCo-SGD to large-scale and high-precision distributed training tasks.
Figure 5: Training time comparison of different methods when the worker node scales. DeCo-SGD demonstrates superior scalability and efficiency compared to other methods across both settings.
5.4 DeCo-SGD under Different Bandwidths and Latency
Table 1 presents the training times of different methods on GPT@Wikitext and ViT@ImageNet under varying network conditions, with all experiments conducted using $4$ worker nodes. The numbers in the table denote the training time (in seconds) required by each method to achieve a target perplexity ($22$ for the GPT model) or accuracy ($80$% for the ViT model). The results clearly demonstrate that under all network conditions, DeCo-SGD consistently outperforms other methods, achieving the lowest training times. This highlights its robustness against bandwidth and latency variations as well as its superior communication efficiency in challenging network environments. For instance, DeCo-SGD achieves up to 5.02×5.02$\times$5.02 × and 1.30×1.30$\times$1.30 × speed-ups over D-SGD and CocktailSGD, respectively.
Table 1: Comparison of training time (s) under different methods and network conditions. $a$ is the average bandwidth and $b$ is the end-to-end latency. Detailed numbers are in the Appendix.
Model @DataSet a(Gbps), b(s) D-SGD Accordion DGA CocktailSGD DeCo-SGD
GPT @Wikitext $0.1,0.1$ $6396.95\pm 32.36$ $1955.28\pm 13.59$ $5506.78\pm 26.77$ $1525.21\pm 12.91$ $\mathbf{1306.29}\pm\mathbf{10.29}$
$0.5,0.1$ $1571.37\pm 11.13$ $829.06\pm 6.65$ $1346.34\pm 10.75$ $712.93\pm 5.38$ $\mathbf{642.79}\pm\mathbf{5.74}$
$0.1,1.0$ $7232.88\pm 35.62$ $2510.26\pm 22.73$ $5764.60\pm 29.69$ $1683.11\pm 14.66$ $\mathbf{1440.59}\pm\mathbf{11.87}$
$0.5,1.0$ $2363.73\pm 13.87$ $1325.12\pm 10.90$ $1562.93\pm 11.89$ $1132.63\pm 9.91$ $\mathbf{997.25}\pm\mathbf{7.92}$
ViT @ImageNet $0.1,0.1$ $1373.84\pm 10.58$ $440.14\pm 3.28$ $1183.68\pm 7.85$ $364.77\pm 1.02$ $\mathbf{283.06}\pm\mathbf{1.80}$
$0.5,0.1$ $446.20\pm 3.39$ $274.60\pm 2.21$ $385.60\pm 1.57$ $219.39\pm 0.98$ $\mathbf{169.13}\pm\mathbf{0.47}$
$0.1,1.0$ $1530.57\pm 9.89$ $604.67\pm 2.79$ $1287.03\pm 10.70$ $505.91\pm 2.56$ $\mathbf{397.16}\pm\mathbf{1.39}$
$0.5,1.0$ $631.15\pm 3.33$ $373.79\pm 1.92$ $444.38\pm 2.18$ $278.57\pm 1.35$ $\mathbf{235.95}\pm\mathbf{0.88}$
6 Related Work
Gradient Compression Strategy. Gradient compression methods are generally categorized into quantization alistarh2017qsgd ; bernstein2018signsgd ; seide20141 , sparsity aji2017sparse ; lin2017deep ; stich2018sparsified , hybrid approaches dryden2016communication ; jiang2018sketchml , and low-rank strategies vogels2019powersgd . Traditional schemes often rely on fixed compression parameters, which limit their adaptability to varying network and training conditions. To address this, recent works have introduced dynamic and adaptive strategies. DC2 abdelmoniem2021dc2 adjusts communication based on real-time network dynamics, while DAGC lu2023dagc assigns compression aggressiveness based on each worker’s data volume. However, neither of them considers end-to-end latency. FedCG jiang2023heterogeneity selects representative clients and adapts compression ratios according to their heterogeneity and capacity changes. L-GreCo markov2024greco formulates the per-layer compression allocation as a constrained optimization problem solved via dynamic programming to balance model size and error. However, neither of the above two methods takes into account the dynamic changes of the network environment.
Asynchronism. FedBuff nguyen2022federated uses a method of buffering asynchronous aggregation to improve communication efficiency and model performance. However, the buffering mechanism may increase memory overhead and delay the incorporation of fresh gradients. Mishchenko et al.mishchenko2022asynchronous rigorously proved theoretically and validated empirically that asynchronous SGD can achieve convergence comparable to mini-batch SGD, even in the presence of arbitrary gradient delays. CO2 sun2024co2 introduces local update and single-step asynchronous communication mechanisms to achieve complete overlap between communication and computing, but it assumes homogeneous device capabilities. To further address heterogeneity in node speed, Shadowheart SGD tyurinshadowheart dynamically adjusts the local update frequency and communication frequency of each node based on the different computing and communication speeds of each working node, thereby achieving the optimal time complexity in an asynchronous setting.
7 Conclusion
In this work, we address the challenge of efficient distributed training in high-latency, low-bandwidth and varying bandwidth networks, where D-SGD suffers significant throughput degradation. While gradient compression and delayed aggregation can mitigate bandwidth and latency issues, respectively, they introduce a complex trade-off among compression ratio, staleness, and model accuracy. To achieve the optimal balance during the whole training, we mathematically model it from two perspectives: time-to-iteration and iteration-to-accuracy. For the former, we derive the threshold compression ratio that maximizes computation-communication overlap under the given staleness and network conditions. For the latter, we introduce NVS to derive the convergence rate of DD-EF-SGD. We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression, filling the critical theoretical gap in DD-EF-SGD. Building on these insights, we propose DeCo-SGD, an adaptive optimizer that dynamically adjusts compression and staleness based on network conditions. Our design achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over the SOTA method. The code will be released upon publication.
Appendix A Notation List
Table 2: Notation list.
Notation Description
$n$ number of workers
$\mathbf{T}$ total number of iterations
$f_{i}(\cdot)$ the local loss function of worker $i$
$f(\cdot)$ the global loss function, i.e., $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$
$\delta$ compression ratio ($0<\delta\leq 1$)
$\mathbf{C}_{\delta}(\cdot)$ the sparsification compressor with compression ratio $\delta$
$\mathbf{x}^{i}_{t}$ the local model parameter in the $t$-th iteration of the $i$-th worker
$\mathbf{x}_{t}$ the global model parameter in the $t$-th iteration
$\mathbf{g}^{i}_{t}$ the stochastic gradient of $\textbf{x}_{t}$ of the worker $i$
$\mathbf{g}_{t}$ the average of $\mathbf{g}^{i}_{t}$, i.e., $\mathbf{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}^{i}_{t}$
$\gamma$ the stepsize
$\tau$ the delayed staleness
$\mathbf{e}_{t}^{i}$ the local error term of the $i$-th worker at the $t$-th iteration
$L$ L-smoothness
$\mu$ $\mu$-strongly convexity if $f_{i}$ is strongly convex
$\zeta_{i}$ the data heterogenority of worker $i$
$\zeta$ the global data heterogenority, i.e., $\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2}$
$\xi^{i}$ the stochastic gradient noise of worker $i$
$\sigma^{2}$ the upper variance bound of the stochastic gradient noise
$a$ the bandwidth (bits/s)
$b$ the network end-to-end latency (s)
$S_{g}$ the size of the gradient (bits)
$T_{comp}$ computation time per iteration (s)
$T_{avg}$ average end-to-end training time per iteration (s)
$T_{sum}$ total end-to-end training time (s)
$TS_{k}$ End time of the k-th computation(s)
$TM_{k}$ End time of the k-th transmission(s)
$TC_{k}$ End time of the k-th communication(s)
Appendix B Detailed proof
B.1 Technical Results
We list lemmas derived from other works here to help us complete the whole proof. Detailed proof of these lemmas can be found from the reference and we do not write here.
Lemma 1.
If $c_{1},c_{2}\in\mathbb{R}^{d}$ then the Jensen’s inequality is: For all $\rho>0$, we have
$\lVert c_{1}+c_{2}\rVert^{2}\leq(1+\rho)\lVert a\rVert^{2}+(1+\rho^{-1})\lVert b\rVert^{2}.$ (12)
This can be written as:
$\langle c_{1},c_{2}\rangle^{2}\leq\rho\lVert a\rVert^{2}+\rho^{-1}\lVert b\rVert^{2}.$ (13)
Lemma 2 (The nature of Top-$k$).
By definition, the Top-$k$ compressor $\mathbf{C}_{\delta}$ is a mapping that has the property $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$:
$\mathbb{E}_{\mathbf{C}_{\delta}}\lVert\mathbf{C}_{\delta}(\mathbf{x})-\mathbf{x}\rVert^{2}\leq(1-\delta)\lVert\mathbf{x}\rVert^{2}.$ (14)
Lemma 3 (Lemma 27 of the work [19]).
Let $(r_{t})_{t\geq 0}$ and $(s_{t})_{t\geq 0}$ be sequences of positive numbers satisfying
$r_{t+1}\leq r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
for some positive constants $B>0$, $C,D\geq 0$ and step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$. Then there exists a constant stepsize $\gamma\leq\frac{1}{E}$ such that
$\frac{B}{T+1}\sum_{t=0}^{T}s_{t}\leq\frac{Er_{0}}{T+1}+2D^{1/3}\left(\frac{r_{0}}{T+1}\right)^{2/3}+2\left(\frac{Cr_{0}}{T+1}\right)^{1/2}.$ (15)
Remark 5.
To ensure that the right hand side in Eq. 15 is less than $\epsilon>0$,
$T=\mathcal{O}\left(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon}\right)\cdot r_{0}$
steps are sufficient.
Lemma 4 (Lemma 25 of the work [19]).
Let $(r_{t})_{t\geq 0}$ and $(s_{t})_{t\geq 0}$ be sequences of positive numbers satisfying
$r_{t+1}\leq(1-\min\{\gamma A,F\})r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
for some positive constants $A,B>0$, $C,D\geq 0$, and for constant step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$, and for parameter $0<F\leq 1$. Then there exists a constant step-size $\gamma\leq\frac{1}{E}$ such that
$\frac{B}{W_{T}}\sum_{t=0}^{T}w_{t}s_{t}+\min\left\{A,\frac{F}{\gamma}\right\}r_{T+1}$ $\leq r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}(T+1)\right]$
$\quad+\frac{2C\ln\tau}{A(T+1)}+\frac{D\ln^{2}\tau}{A^{2}(T+1)^{2}}$
for $w_{t}:=(1-\min\{\gamma A,F\})^{-(t+1)}$, $W_{T}:=\sum_{t=0}^{T}w_{t}$ and
$\tau=\max\left\{\exp[1],\min\left\{\frac{A^{2}r_{0}(T+1)^{2}}{C},\frac{A^{3}r_{0}(T+1)^{3}}{D}\right\}\right\}$
Remark 6.
Lemma  4 establishes a bound of the order
$\tilde{\mathcal{O}}\left(r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}T\right]+\frac{C}{AT}+\frac{D}{A^{2}T^{2}}\right),$
that decreases with $T$. To ensure that this expression is less than $\epsilon$,
$T=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}\log\frac{1}{\epsilon}+\frac{E}{A}\log\frac{1}{\epsilon}\right)=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}+\frac{E}{A}\right)$
steps are sufficient.
B.2 Proof of Theorem 1
Lemma 5.
Let $\left\{\mathbf{x}_{t},\tilde{\mathbf{x}}_{t},\hat{\mathbf{x}}_{t},B_{t},\tilde{B}_{t},\mathbf{v}_{t},\tilde{\mathbf{v}}_{t}\right\}_{t\geq 0}$ be defined as in DD-EF-SGD using NVS, with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. If $\gamma\leq\frac{1}{4L}$, then we have
$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+\gamma L^{2}\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+2\gamma L^{2}\mathbb{E}\left\|B_{t}\right\|^{2}+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
Proof. We begin with the definition of $\hat{\mathbf{x}}_{t+1}$ and have
$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)$ $=\mathbb{E}f\left(\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right)$ (16)
$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\frac{L}{2}\mathbb{E}\left\|\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right\|^{2}$
$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
The first line is due to $\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$. The second line is due to Eq. 1 and the third line is due to Eq. 3. We estimate $A_{1}$ as
$A_{1}=\gamma\mathbb{E}\langle-\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\rangle$ (17)
$=\gamma(\mathbb{E}\left\langle\nabla f\left(\tilde{\mathbf{x}}_{t}\right)-\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right)-\nabla f\left(\tilde{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle-\mathbb{E}\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{g}_{t}\rangle)$
$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{c_{2}\gamma L^{2}}{2}\mathbb{E}\lVert B_{t}\rVert^{2}+\frac{\gamma}{2c_{2}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}.$
The last line is due to Jensen’s inequality (i.e., Eq. 13). We let $c_{1}=2$, $c_{2}=4$. By bringing Eq. LABEL:A-1 into Eq. 16, we complete the proof.
Lemma 6.
Define $\{\tilde{B}_{t}\}_{t\geq 0}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{4L\sqrt{\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{16L^{2}}\sum_{t=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{\tau\gamma^{2}(1+T)\sigma^{2}}{n}.$
Proof. This lemma follows from the Lemma 20 in the previous work [14].
Lemma 7.
Define $\{B_{t}\}_{t\geq 0}$ as in Eq. 3.2. Then by using the relative compressor, with $\gamma\leq\frac{1}{4L\sqrt{\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert B_{t}\rVert^{2}\leq\frac{1}{4L^{2}}\mathbb{E}\sum_{t=0}^{T}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi(T+1)(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Proof.
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\frac{(1-\delta)(1+\rho)}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}$
$+\gamma^{2}(1-\delta)((1+\rho^{-1})\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+(1+\rho^{-1})\zeta^{2}+\sigma^{2}).$
This inequality is due to Lemma 2, Eq. 4, 3,12.
By unrolling the recurrence and let $\rho=\frac{2-\delta}{\delta}$, we can derive the upper bound:
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we have
$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]\leq\frac{1}{4L^{2}}\sum_{i=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2\gamma^{2}(1-\delta)(1+T)}{\delta(1-\frac{\delta}{2})^{\tau}}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
Theorem 1 (Non-convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}L}{n\epsilon^{2}}+\frac{L\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{L}{\epsilon}+\frac{L\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}L}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
iterations of DD-EF-SGD, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total iterations, chosen at random uniformly.
Proof. Let $G_{t}:=\sum_{i=0}^{t}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{i}\right)\right\|^{2}$. According to above lemmas (Lemma 5, 6, 7), we have
$\frac{\gamma}{2}G_{T}\leq\mathbb{E}f(\hat{\mathbf{x}}_{0})-\mathbb{E}f(\hat{\mathbf{x}}_{T+1})+\gamma L^{2}(\sum_{i=0}^{T}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}\lVert B_{t}\rVert^{2})+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}$
$\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma}{4}G_{T}+\frac{\gamma}{8}G_{T}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+2\gamma^{3}L^{2}(T+1)(2\phi\frac{2}{\delta}\zeta^{2}+2\phi\sigma^{2}+\frac{2\tau\sigma^{2}}{n}).$
Divide both sides of the inequality by $4\gamma(1+T)$, we have:
$\frac{1}{8(1+T)}G_{T}\leq\frac{\mathbb{E}f(\mathbf{x}_{0})-f^{*}}{\gamma(T+1)}+\gamma\frac{L\sigma^{2}}{2n}+4\gamma^{2}L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n}).$
According to Lemma 3, we let $B=1/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n})$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$. To ensure the right hand side of the above inequality is less than $\epsilon>0$, $T=\mathcal{O}(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon})r_{0}$ iterations are sufficient, which completes the proof.
B.3 Proof of Theorem 2
For further discussion, let’s define $F_{t}:=\mathbb{E}f(\mathbf{x}_{t})-f^{*}$ and $X_{t}:=\mathbb{E}\lVert\hat{\mathbf{x}}_{t}-\mathbf{x}^{*}\rVert^{2}$.
Lemma 8.
Let $\left\{\mathbf{x}_{t},\tilde{\mathbf{x}}_{t},\hat{\mathbf{x}}_{t},B_{t},\tilde{B}_{t},\mathbf{v}_{t},\tilde{\mathbf{v}}_{t}\right\}_{t\geq 0}$ be defined as in DD-EF-SGD using NVS, with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-strongly convex. If $\gamma\leq\frac{1}{4L}$, then we have
$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+\gamma^{2}\sigma^{2}+6\gamma L(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
Proof.
$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-x^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
We estimate the term $A_{2}$ as
$A_{2}$ $=-2\gamma(\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\mathbf{x}^{*}\right\rangle-\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\tilde{\mathbf{x}}_{t}-\hat{\mathbf{x}}_{t}\right\rangle-\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\right\rangle)$
$\leq-\gamma\mu\mathbb{E}\lVert\mathbf{x}_{t}-\mathbf{x}^{*}\rVert^{2}-2\gamma F_{t}+\gamma F_{t}+4\gamma L\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+4\gamma L\mathbb{E}\lVert B_{t}\rVert^{2}$
$\leq-\frac{\gamma\mu}{2}X_{t}-\gamma F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
The second line is due to the reason that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{4L},\mu\leq L$, we have
$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}+(2L\gamma^{2}-\gamma)F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})$
$\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}f_{t}+6L\gamma(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
This completes the proof.
Lemma 9.
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $\tilde{B}_{t}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{8\sqrt{L\tau}}$, we have
$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n},$
where $W_{t}=\sum_{i=1}^{n}w_{i}$.
Proof. This lemma follows from the Lemma 20 in the previous work [14].
Lemma 10.
Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $B_{t}$ as in Eq. 3.2. Then by using the relative compressor, with $\gamma\leq\frac{1}{16\sqrt{L\phi}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$ we have
$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
where $W_{t}=\sum_{i=1}^{n}w_{i}$.
proof. Denote $E_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$, and we have:
$w_{t}E_{t}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}w_{i}(1+\frac{\delta}{4})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2})$
$+\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}$
$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{i=0}^{t}(1-\frac{\delta}{4})^{t-i}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}]$
$\sum_{t=0}^{T}w_{t}E_{t}$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{j=0}^{\infty}(1-\frac{\delta}{4})^{j}\sum_{i=0}^{T}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{4}{\delta}\sum_{i=0}^{T}w_{i}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})].$
Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we complete the proof.
Theorem 2 (Convex convergence rate of DD-EF-SGD).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex. Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$
iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
Proof. By substituting Lemma 5, 6 into Lemma 4, we can get
$X_{t}\leq w_{T}X_{0}-\frac{\gamma}{2}\sum_{i=0}^{T}w_{t}F_{t}+\frac{W_{T}\gamma^{2}\sigma^{2}}{n}+6\gamma L(\sum_{i=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2})$
$\leq w_{T}X_{0}-\frac{1}{8}\sum_{i=0}^{T}w_{t}F_{t}+W_{T}\gamma^{2}\sigma^{2}+6LW_{T}\gamma^{3}(\frac{\tau\sigma^{2}}{n}+\frac{4\phi\zeta^{2}}{\delta}+2\phi\sigma^{2}).$
According to Lemma 4, we complete the proof.
B.4 Proof of Theorem 3
Theorem 3 (Estimation of $T_{avg}$ in DD-EF-SGD).
Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed. Then the average iteration time $T_{avg}:=\frac{TC_{t}}{t}$ satisfies the following approximation:
$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1},\frac{\delta S_{g}}{a},T_{\text{comp}}\right\}},$ (18)
where the approximation error is upper bounded by $O(\frac{1}{t})$.
proof. Based on the assumption, the iterative formula of $TS_{k}$ derived from the relationship between computation and communication can be written as :
$\left\{\begin{matrix}TC_{k}=TM_{k}+b\\ TS_{k+1}=T_{\text{comp}}+max\left\{TC_{k-\tau},TS_{k}\right\}\\ TM_{k+1}=\frac{\delta S_{g}}{a}+max\left\{TM_{k},TS_{k+1}\right\}\end{matrix}\right.\quad,k=0,1,2...\quad.$ (19)
with initial values $TS_{0}=TM_{0}=0$ and $TC_{k}=0$ for all $k\leq 0$. The analysis is divided into four cases, each corresponding to a specific regime determined by the relative magnitudes of $T_{\text{comp}}$, $\frac{\delta S_{g}}{a}$, and $\tau$.
Case 1: Computation-Dominated Regime
Consider the regime where $T_{\text{comp}}>\frac{\delta S_{g}}{a}$ and $\tau T_{\text{comp}}>\frac{\delta S_{g}}{a}+b$. We claim that:
$\forall k\in\mathbb{N}^{+},\quad TS_{k}=kT_{\text{comp}}.$
To verify this, we define the inductive hypothesis:
$P(k):\quad\left\{\begin{aligned} TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\forall i=1,\ldots,k.$
The base case $P(1)$ can be directly verified from the recurrence relations. Now assume $P(k)$ holds for some $k\in\mathbb{N}^{+}$. We analyze the $(k+1)$-th step:
  • If $k\leq\tau$, then $TC_{k-\tau}=0\leq TS_{k}$;
  • Otherwise, by the inductive assumption, we have:
    $TC_{k-\tau}=(k-\tau)T_{\text{comp}}+\frac{\delta S_{g}}{a}+b\leq kT_{\text{comp}}=TS_{k}.$
Thus, it follows that:
$TS_{k+1}=T_{\text{comp}}+\max\{TC_{k-\tau},TS_{k}\}=T_{\text{comp}}+TS_{k}=(k+1)T_{\text{comp}}.$
Moreover,
$TM_{k+1}=\frac{\delta S_{g}}{a}+\max\{TM_{k},TS_{k+1}\}=\frac{\delta S_{g}}{a}+TS_{k+1},$
since $TM_{k}\leq TS_{k+1}$ by the inductive assumption. Therefore, $P(k+1)$ holds, completing the induction.
It follows that:
$TC_{t}=TM_{t}+b=\frac{\delta S_{g}}{a}+tT_{\text{comp}}+b,$
and thus the approximation error satisfies:
$\left|TC_{t}-tT_{\text{comp}}\right|=\frac{\delta S_{g}}{a}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
Case 2: Communication-Dominated Regime
Now consider the case where $\frac{\delta S_{g}}{a}>T_{\text{comp}}$ and $\tau\cdot\frac{\delta S_{g}}{a}>T_{\text{comp}}+b$. We claim that there exists $N_{1}\in\mathbb{N}^{+}$ such that:
$\forall k>N_{1},\quad TS_{k+1}-TS_{k}=\frac{\delta S_{g}}{a}.$
To establish this, we define the inductive statement:
$P(k):\quad\left\{\begin{aligned} TM_{i}&=T_{\text{comp}}+i\cdot\frac{\delta S_{g}}{a},\\ TM_{i}&\geq TS_{i+1},\end{aligned}\right.\quad\forall i=1,\ldots,k.$
For $i=1,\ldots,\tau+1$, direct computation gives:
$TS_{i}=iT_{\text{comp}},\quad TM_{i}=T_{\text{comp}}+i\cdot\frac{\delta S_{g}}{a},$
hence $P(\tau)$ holds. Assume $P(k)$ is true for some $k\geq\tau$.
Then:
$TM_{k+1}=\frac{\delta S_{g}}{a}+TM_{k}=T_{\text{comp}}+(k+1)\cdot\frac{\delta S_{g}}{a}.$
For the next computation step:
$TC_{k+1-\tau}+T_{\text{comp}}$ $=(k+1-\tau)\cdot\frac{\delta S_{g}}{a}+2T_{\text{comp}}+b$
$\leq(k+1)\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}\leq TM_{k+1},$
$TS_{k+1}+T_{\text{comp}}\leq TM_{k}+T_{\text{comp}}\leq TM_{k+1}.$
Therefore,
$TS_{k+2}=T_{\text{comp}}+\max\left\{TC_{k+1-\tau},TS_{k+1}\right\}\leq TM_{k+1},$
which implies $P(k+1)$ holds. By induction, $P(k)$ is valid for all $k\geq\tau$.
Consequently, for all sufficiently large $k$, the recurrence stabilizes and we obtain:
$TS_{k}=T_{\text{comp}}+TC_{k-\tau}=2T_{\text{comp}}+b+(k-\tau)\cdot\frac{\delta S_{g}}{a},$
with:
$\left|TC_{t}-t\cdot\frac{\delta S_{g}}{a}\right|=T_{\text{comp}}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
Finally, we verify the existence of $N_{1}>\tau$ such that $TC_{N_{1}-\tau}\geq TS_{N_{1}}$. If such $N_{1}$ does not exist, then both $\{TC_{k-\tau}\}$ and $\{TS_{k}\}$ would form arithmetic sequences with conflicting common differences, leading to a contradiction.
Case 3: Periodic Structure under Intermediate Delays
Consider the case where the computation time satisfies $T_{\text{comp}}>\frac{\delta S_{g}}{a}$ while the aggregated period satisfies $\tau T_{\text{comp}}\leq\frac{\delta S_{g}}{a}+b$. Under this condition, we can establish the existence of a positive integer $N$ such that the sequence $\left\{TS_{k+1}-TS_{k}\right\}_{k=N}^{T-1}$ becomes $(\tau+1)$-periodic. Specifically, for all $k\geq N$ and any $d\in\{1,\dots,\tau+1\}$, we have:
$TS_{(k+1)(\tau+1)+d}-TS_{k(\tau+1)+d}=T_{\text{comp}}+b+\frac{\delta S_{g}}{a}.$
This periodic property can be verified by explicitly analyzing the first $(2\tau+2)$ terms of the sequence $\{(TS_{k},TM_{k})\}$. We have:
{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\text{for }i=1,\dots,\tau+1;$ (20)
{$TS_{i}&=\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\\ TM_{i}&=2\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\end{aligned}\right.\quad\text{for }i=\tau+2,\dots,2\tau+2.$
Let us define the index group:
$A_{k}$ :=$:=\left\{(TM_{k(\tau+1)+d},TS_{k(\tau+1)+d})\mid d=1,\dots,\tau+1\right\},$
$A_{k}+a$ :=$:=\left\{(TM_{k(\tau+1)+d}+a,TS_{k(\tau+1)+d}+a)\mid d=1,\dots,\tau+1\right\},\quad\forall a\in\mathbb{R}.$
By rewriting the iterative update rule in the compact form $A_{k}=F(A_{k-1})$, it follows from translation invariance that:
$A_{k}+a=F(A_{k-1}+a).$
Now define the shifted sequence $A_{k}^{\prime}:=A_{k+1}-\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)$. Using Eq. (20), we obtain:
$A_{0}^{\prime}$ $=A_{1}-\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)=A_{0},$
$A_{1}^{\prime}$ $=F(A_{0}^{\prime})=A_{1}.$
Therefore, we conclude:
$A_{2}$ $=A_{1}^{\prime}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)=A_{1}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right),$
and recursively,
$A_{k+1}=A_{k}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right),\quad\forall k\in\mathbb{N}^{*}.$
Let $t=m(\tau+1)+d$ for $d=1,\dots,\tau+1$, then the cumulative time $TC_{t}$ satisfies:
$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+dT_{\text{comp}}+\frac{\delta S_{g}}{a}+b,$
$\left|TC_{t}-t\left(\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)\right|$ $=d\left(T_{\text{comp}}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+\frac{\delta S_{g}}{a}+b$
$\leq\tau T_{\text{comp}}\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
This confirms that $TC_{t}$ deviates from its average rate by at most a bounded constant, further supporting the regularity and predictability of the update pattern in this intermediate-delay regime.
Case 4: Periodic Structure under Computation-Dominated Regime
In this case, we consider the regime where $T_{\text{comp}}<\frac{\delta S_{g}}{a}$ and $\tau\frac{\delta S_{g}}{a}\leq T_{\text{comp}}+b$. Under these conditions, we can establish that there exists a positive integer $N$ such that the update interval sequence $\left\{TS_{k+1}-TS_{k}\right\}_{k=N}^{T-1}$ exhibits $(\tau+1)$-periodicity. In particular, for all $k\geq N$ and for all $d\in\{1,\dots,\tau+1\}$, the following recurrence holds:
$TS_{(k+1)(\tau+1)+d}-TS_{k(\tau+1)+d}=T_{\text{comp}}+b+\frac{\delta S_{g}}{a}.$
Analogously to Case 3, we examine the first $(3\tau+3)$ elements of the sequence $\left\{(TS_{k},TM_{k})\right\}$ to validate the periodic structure:
{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=i\frac{\delta S_{g}}{a}+T_{\text{comp}},\end{aligned}\right.\quad$ $i=1,\dots,\tau+1;$
{$TS_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-2)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-1)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=\tau+2,\dots,2\tau+2;$
{$TS_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-3)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-2)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=2\tau+3,\dots,3\tau+3.$
We define the grouping $A_{k}$ as in Case 3. From the recurrence observed above, we find that:
$A_{2}=A_{1}+\left(T_{\text{comp}}+\frac{\delta S_{g}}{a}+b\right),$
and by translation invariance of the update dynamics, it follows that:
$A_{k+2}=A_{k+1}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right),\quad\forall k\in\mathbb{N}^{*}.$
Thus, the $(\tau+1)$-periodicity is established, completing the proof for Case 4.
Now, let $t=m(\tau+1)+d$ with $d\in\{1,\dots,\tau+1\}$. The cumulative computation time $TC_{t}$ satisfies:
$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+d\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}+b,$
$\left|TC_{t}-t\cdot\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right|$ $=d\left(\frac{\delta S_{g}}{a}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+T_{\text{comp}}+b$
$\leq\tau\cdot\frac{\delta S_{g}}{a}\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
Summary of All Cases and Average Iteration Time
By integrating the results from all four cases, we obtain a unified expression for the asymptotic average iteration time:
$T^{\prime}_{\text{avg}}$ =$=\begin{cases}\dfrac{T_{\text{comp}}+\frac{\delta S_{g}}{a}+b}{\tau+1},&\tau<\tau_{\text{thres}},\\ \max\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},&\tau\geq\tau_{\text{thres}},\end{cases}\quad\text{with}\quad\left|TC_{t}-tT^{\prime}_{\text{avg}}\right|\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},$
where $T^{\prime}_{\text{avg}}:=\lim\limits_{k\to\infty}\frac{TS_{k}}{k}$ approximates the true average iteration time $T_{\text{avg}}$, and
$\tau_{\text{thres}}=\min\left\{\frac{\frac{\delta S_{g}}{a}+b}{T_{\text{comp}}},\frac{T_{\text{comp}}+b}{\frac{\delta S_{g}}{a}}\right\}.$
As a final conclusion, we estimate the average iteration time by:
$T_{\text{avg}}\approx T^{\prime}_{\text{avg}}=\max\left\{\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1},\frac{\delta S_{g}}{a},T_{\text{comp}}\right\},$
with the absolute error bounded by:
$\left|T_{\text{avg}}-T^{\prime}_{\text{avg}}\right|\leq\frac{b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}}{t}=O\left(\frac{1}{t}\right).$
Appendix C Addendum to Evaluation Experiments
C.1 Experimental Environment
The experiments are conducted on an Ubuntu $24.04$ LTS server with an Intel Xeon Gold $6230$ CPU and $3$ Nvidia A40404040 GPUs, each with $48$GB memory. The Python version is $3.10.16$, and other used libraries are all based on the Python version. We use PyTorch $2.5.1$ with CUDA $12.4$ as the ML toolkit. We use gloo as the communication backend.
C.2 Detailed Experimental Settings
Experiment tasks: The experimental setup encompasses three task types spanning computer vision (CV) and natural language processing (NLP), utilizing both traditional (CNN [29]) and state-of-the-art transformer-based models (ViT [30] and GPT [31]). The CNN architecture includes two convolutional layers followed by two fully connected layers, used widely in previous works [9, 20]. The ViT base model features approximately 86 million parameters, while the smallest GPT variant contains 124 million parameters.
To demonstrate the superiority of our design, we evaluate four representative tasks: FashionMNIST [32] and CIFAR-$10$ with CNN (denoted as CNN@FMNIST and CNN@CIFAR-$10$), ImageNet [33] using ViT (denoted as ViT@ImageNet) and Wikitext [34] on GPT (denoted as GPT@Wikitext). For CNN@FMNIST and CNN@CIFAR-$10$, the learning rate is $0.01$, and the batch size is $32$. For ViT@ImageNet and GPT@Wikitext, the learning rates are $0.01$ and $0.1$, respectively, with batch sizes of $40$ and $5$.
Baselines: We compare D-SGD, Accordion, DGA, and vanilla DD-EF-SGD as baselines. D-SGD is the case $\tau=0,\delta=1$, and While it does not suffer from accuracy degradation, it converges the slowest in high-latency and low-bandwidth scenarios when measured in terms of end-to-end time. Accordion [35] is the SOTA adaptive gradient compression algorithm, which sets small $\delta$ in the critical iterations and sets large $\delta$ if not [9]. The compressor we use in all baselines is Top-$k$, which transmits the largest $k$ elements of gradients. The reason why we select Top-$k$ is its strong empirical performance, widespread adoption in DML, and straightforward implementation. DGA is a state-of-the-art algorithm originally designed for Federated Learning [18], which first introduces delay staleness to reduce the negative effect of the high latency on Federated Learning. We set $K=1$ in DGA to close the infrequent communication. CocktailSGD [21] uses the fixed $\tau$ and $\delta$. The hyperparameters are chosen by DeCo-SGD with $E=\infty$.
Network environment: We simulated challenging network environments (static high latency $\geq$ $100$ms, dynamic low bandwidth, average bandwidth $\leq$ $1$Gbps), where the dynamic volatility of the network can be referred to the following figure, which gives the network environment in the given environment (Section 5.2), and we discuss the superiority of the DeCo-SGD algorithm in different network conditions (Section 5.4).
C.3 Network Conditions of Section 5.2
As shown in Fig. LABEL:fig:network, the network conditions exhibit dynamic bandwidth fluctuations across all tasks (CNN@FMNIST, CNN@CIFAR-10, ViT@ImageNet, and GPT@Wikitext), and the compression ratio $\delta$ is adaptively adjusted over time under a fixed 200 ms transmission delay with $\tau=2$.
C.4 Detailed Curves of Section 5.3 to Verify the Scalability of DeCo-SGD
The detail of Fig. 5 in the main text of the paper corresponds to Fig. 9 in the Appendix.
(a) ViT@ImageNet $n=4$
(b) ViT@ImageNet $n=8$
(c) ViT@ImageNet $n=16$
(d) ViT@ImageNet $n=32$
(e) GPT@Wikitext $n=4$
(f) GPT@Wikitext $n=8$
(g) GPT@Wikitext $n=16$
(h) GPT@Wikitext $n=32$
Figure 9: Convergence performance across different numbers of worker nodes ($n=4,8,16,32$) for five distributed optimization methods. Top row: Accuracy over time for ViT@ImageNet. Bottom row: Perplexity over time for GPT@Wikitext.
C.5 Detailed Results of Section 5.4
Fig. 10 in the Appendix provides a visual counterpart to the data presented in Table 1 in the main text. Table 3 extends more details in Table 1
Table 3: Comparison of training time (s) under different methods and network conditions. $a$ is the average bandwidth and $b$ is the end-to-end latency. $\tau^{*}$ and $\delta^{*}$ are caculated by DeCo-SGD. The numbers in parentheses represent the speedup compared to DeCo-SGD.
Model @DataSet a(Gbps), b(s) $\tau^{*}$, $\delta^{*}$ D-SGD Accordion DGA CocktailSGD DeCo-SGD
GPT @Wikitext $0.1,0.1$ $2,0.02$ $6396.95$ (4.90×4.90$\times$4.90 ×) $1955.28$ (1.50×1.50$\times$1.50 ×) $5506.78$ (4.22×4.$2$) $1525.21$ (1.17×1.17$\times$1.17 ×) $\mathbf{1306.29}$
$0.5,0.1$ $2,0.10$ $1571.37$ (2.44×2.$4$) $829.06$ (1.29×1.29$\times$1.29 ×) $1346.34$ (2.09×2.09$\times$2.09 ×) $712.93$ ($1\times1$) $\mathbf{642.79}$
$0.1,1.0$ $3,0.02$ $7232.88$ (5.02×5.02$\times$5.02 ×) $2510.26$ (1.74×1.74$\times$1.74 ×) $5764.60$ (4.00×4.$0$) $1683.11$ (1.17×1.17$\times$1.17 ×) $\mathbf{1440.59}$
$0.5,1.0$ $3,0.10$ $2363.73$ (2.37×2.37$\times$2.37 ×) $1325.12$ (1.33×1.$3$) $1562.93$ (1.57×1.57$\times$1.57 ×) $1132.63$ (1.14×1.14$\times$1.14 ×) $\mathbf{997.25}$
ViT @ImageNet $0.1,0.1$ $2,0.03$ $1373.84$ (4.85×4.85$\times$4.85 ×) $440.14$ (1.55×1.$5$) $1183.68$ (4.18×4.18$\times$4.18 ×) $364.77$ (1.29×1.29$\times$1.29 ×) $\mathbf{283.06}$
$0.5,0.1$ $2,0.14$ $446.20$ (2.64×2.64$\times$2.64 ×) $274.60$ (1.62×1.62$\times$1.62 ×) $385.60$ (2.28×2.28$\times$2.28 ×) $219.39$ (1.30×1.30$\times$1.30 ×) $\mathbf{169.13}$
$0.1,1.0$ $3,0.03$ $1530.57$ (3.85×3.85$\times$3.85 ×) $604.67$ (1.52×1.52$\times$1.52 ×) $1287.03$ (3.24×3.24$\times$3.24 ×) $505.91$ (1.27×1.27$\times$1.27 ×) $\mathbf{397.16}$
$0.5,1.0$ $3,0.14$ $631.15$ (2.67×2.67$\times$2.67 ×) $373.79$ (1.58×1.58$\times$1.58 ×) $444.38$ (1.88×1.$8$) $278.57$ (1.18×1.18$\times$1.18 ×) $\mathbf{235.95}$
(a) ViT@ImageNet bandwidth=$100$Mbps latency=$100$ms
(b) ViT@ImageNet bandwidth=$100$Mbps latency=$1000$ms
(c) ViT@ImageNet bandwidth=$500$Mbps latency=$100$ms
(d) ViT@ImageNet bandwidth=$500$Mbps latency=$1000$ms
(e) GPT@Wikitext bandwidth=$100$Mbps latency=$100$ms
(f) GPT@Wikitext bandwidth=$100$Mbps latency=$1000$ms
(g) GPT@Wikitext bandwidth=$500$Mbps latency=$100$ms
(h) GPT@Wikitext bandwidth=$500$Mbps latency=$1000$ms
Figure 10: Convergence performance across different numbers of worker nodes ($n=4,8,16,32$) for five distributed optimization methods. Top row: Accuracy over time for ViT@ImageNet. Bottom row: Perplexity over time for GPT@Wikitext.
Appendix D Others
D.1 UsedResources and Licenses
All external assets used in this paper are properly credited, and their licenses and terms of use have been explicitly acknowledged and respected. Specifically, we make use of several open-source resources as follows:
  1. 1.
    ViT model, released under the MIT License.
  2. 2.
    GPT2 model from Hugging Face, released under the MIT License.
  3. 3.
    docker-tc, used to simulate network conditions, released under the MIT License.
All of these assets were used strictly for research purposes in accordance with their respective licenses. We have cited the corresponding authors or repositories in our references where appropriate and ensured full compliance with their terms of use.
D.2 Limitations and Boarder Impact
Limitations: (1) Although DeCo-SGD currently supports only the D-SGD and its variants, we believe that D-SGD is applicable to a wide range of tasks, including NLP, CV and Objective Detection. If D-SGD is properly parameterized, its performance is no worse than popular optimizers in LLM training such as Adam [52]. (2) Device heterogeneity (different bandwidth/latency per node) is not considered, as current theoretical frameworks (based on convex optimization) cannot analyze such scenarios without prohibitive parameter search costs. Nevertheless, our approach already achieves $80\%$ throughput gains over SOTA optimizer [21] in dynamic networks, leaving limited room for further improvement via heterogeneity optimization.
Boarder Impact: Our work establishes theoretical foundations for LLM pre-training in low-bandwidth, high-latency networks. While employing lossy training techniques, we provide theoretical guarantees of training efficiency. This could significantly reduce the infrastructure costs for LLM pre-training, promoting broader participation in AI research and development.

### v2

@@
-DeCo-SGD: Joint Optimization of Delay Staleness and Gradient Compression Ratio for Distributed SGD
+Taming Latency and Bandwidth: A Theoretical Framework and Adaptive Algorithm for Communication-Constrained Training
 Abstract
-Distributed machine learning in high end-to-end latency and low, varying bandwidth network environments undergoes severe throughput degradation.
-Due to its low communication requirements, distributed SGD (D-SGD) remains the mainstream optimizer in such challenging networks, but it still suffers from significant throughput reduction.
-To mitigate these limitations, existing approaches typically employ gradient compression and delayed aggregation to alleviate low bandwidth and high latency, respectively.
-To address both challenges simultaneously, these strategies are often combined, introducing a complex three-way trade-off among compression ratio, staleness (delayed synchronization steps), and model convergence rate.
-To achieve the balance under varying bandwidth conditions, an adaptive policy is required to dynamically adjust these parameters.
-Unfortunately, existing works rely on static heuristic strategies due to the lack of theoretical guidance, which prevents them from achieving this goal.
-This study fills in this theoretical gap by introducing a new theoretical tool, decomposing the joint optimization problem into a traditional convergence rate analysis with multiple analyzable noise terms.
-We are the first to reveal that staleness exponentially amplifies the negative impact of gradient compression on training performance, filling a critical gap in understanding how compressed and delayed gradients affect training.
-Furthermore, by integrating the convergence rate with a network-aware time minimization condition, we propose DeCo-SGD, which dynamically adjusts the compression ratio and staleness based on the real-time network condition and training task.
-DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and static strategy in high-latency and low, varying bandwidth networks, respectively.
+Regional energy caps limit the growth of any single data center used for large-scale model training.
+This single-center training paradigm works when model size remains manageable, but exponential growth in the model size and computational demand challenges it.
+A natural alternative is to distribute training across multiple data centers over wide-area networks.
+This pools distributed resources, but suffers from high latency and low, time-varying bandwidth, sharply reducing throughout.
+Employing jointly gradient compression and delayed aggregation can alleviate communication problems, but introduces a complex three-way trade-off among compression ratio, staleness (delayed synchronization steps), and convergence rate.
+Existing work lacks theoretical guidance and can only propose fixed strategies, insensitive to computation and communication conditions.
+We address this with a new theoretical tool, decomposing the joint optimization problem into a traditional process plus multiple analyzable noise terms.
+Our analysis yields the first convergence rate for this setting and shows that increasing staleness exponentially amplifies the detrimental effect of compression.
+Leveraging these insights, we propose DeCo-SGD, which dynamically selects the compression ratio and staleness based on the real-time communication and computation conditions.
+DeCo-SGD achieves up to $5.07\times$ and $1.37\times$ speed-ups over distributed SGD and static strategy in high-latency and low, varying bandwidth networks, respectively.
 1 Introduction
-Recent breakthroughs in large language models (LLMs) (e.g., GPT-4 gpt-4 , DeepSeek-v3 deepseek ) have been accompanied by an explosive growth in model parameters, leading to unprecedented communication demands for distributed training scalalaws2 ; ScalaLaws1 .
-State-of-the-art (SOTA) distributed training strategies flashattention ; megatron-lm2 are predominantly optimized for high-speed interconnects like PCIe 5.0 and NVLink nvidia , where bandwidth exceeds $10$ Gbps and latency remains under $10$ ms.
-Although these assumptions are reasonable in well-resourced data centers, they often fail to hold in dynamic network environments with high latency and low bandwidth jaghouar2024opendiloco .
-Unfortunately, such unfavorable network conditions are common in distributed training over wide area networks (WAN), such as federated learning dagctmc and cross-geographic distributed learning hsieh2017gaia .
-WAN-based distributed training lowers the barrier for LLM pre-training, empowering individual researchers and small labs while reducing the reliance on costly, specialized networking SIGCOMM24StellaTrain .
-In these challenging dynamic network settings, distributed SGD (D-SGD) and its variants, such as FedAVG fedavg and D-QSGD cui2021slashing , remain the dominant optimization methods due to their communication-friendly nature.
-Despite its low communication requirements, D-SGD still suffers from significant throughput degradation under high-latency and low-bandwidth conditions.
-As illustrated in Fig.
-1, when the bandwidth drops below $2$ Gbps and the latency exceeds $200$ ms, the throughput of D-SGD in training GPT-2 with four nodes reduces to $50\%$ of maximum achievable throughput [^1] The maximum achievable throughput is determined by the compute capacity of the machines, without network constraints..
-{wrapfigure}t0.4 [Uncaptioned image]
-[…]
+The rapid progress of large language models (LLMs) (e.g., GPT-4 gpt-4, DeepSeek-v3 deepseek) has been accompanied by an explosive growth in model parameters.
+This expansion drives a steep rise in the computational resources required for training scalalaws2; ScalaLaws1.
+Training in a single AI computing center, where thousands of GPUs are tightly coupled via high-speed interconnects (e.g., PCIe and NVLink nvidia), is increasingly constrained by regional energy capacity.
+This bottleneck is already evident in major hubs; for instance, data centers in Virginia are projected to consume up to $51\%$ of the state’s electricity by 2030 GEI-Datacenter-25, while utilities in Ohio have imposed moratoriums on new connections due to transmission limits Williams2024.
+As model sizes continue to grow exponentially, the paradigm of training in a hyperscale compute center is approaching its scalability limits.
+This observation motivates a new phase of large-scale training that extends beyond a single compute center or energy domain.
+Distributed training over wide-area networks (WANs) offers a practical and sustainable way to overcome regional power limits and coordinate multiple compute centers globally.
+Recently, some pioneering projects have already demonstrated the feasibility of training over WANs as shown in Fig.
+2.
+Intellect-1 collaboratively trained a 1010B-parameter model on up to $112$ H100100s distributed across five countries and three continents jaghouar2024opendiloco.
+Similarly, DeepLink connected datacenters across two cities in China, more than 1,5001{,}500 km apart, to train a model exceeding 100100B parameters, highlighting the potential of multi-region training over WANs to help alleviate regional power constraints and begin to harness geographically distributed compute at scale deeplink.
+Distributed training over WANs is fundamentally communication-constrained, facing high latency and limited bandwidth.
+In this setting, the standard data-parallel paradigm requires workers to perform local computation and then exchange full gradients with the server.
+Then the server aggregates the gradients, updates the model, and broadcasts the new parameters to workers.
+Among algorithms employing this paradigm, Distributed Stochastic Gradient Descent (D-SGD) and its variants are the most widely adopted, thanks to their clean iterative formulation and practical scalability dagctmc; SIGCOMM24StellaTrain.
+However, under WAN conditions, the need to transmit full gradients leads to severe degradation in end-to-end training throughput (i.e., the number of samples processed per second) for these methods dong25emnlpsurvey.
+Figure 1: Pioneering LLM training over WANs.
+DeepLink connects compute clusters in Shanghai and Ji’nan over 1,5001{,}500 km, with an estimated throughput efficiency of $35\%$.
+Intellect-1 trains across continents, where the corresponding throughput efficiency drops to approximately $15\%$.
+Figure 2: The heatmap of throughput efficiency ($\%$) for D-SGD with four nodes training GPT-2, under different latency and bandwidth conditions.
+Each node is equipped with one A4040 GPU.
+The throughput efficiency at (x, y) is defined as the throughput at (x, y) divided by the maximum achievable throughput of the machines.
+To quantify the impact of network conditions, we conduct some controlled experiments training GPT-2 under varying bandwidth and latency configurations.
+Fig.
+2 shows the resulting throughput efficiency, defined as end-to-end training throughput normalized by the maximum achievable throughput [^1] This is determined by the compute capacity of the machines, without network constraints..
+By mapping real-world parameters to this heatmap, we observe severe degradation: DeepLink (roughly 1,0001{,}000 Mb/s bandwidth and $100$ ms latency amazon) falls into a region with only about $35\%$ of the ideal throughput.
+Similarly, Intellect-1 (roughly $100$ Mb/s bandwidth and $500$ ms latency amazon) lies in a region with throughput efficiency around $15\%$.
+Taken together, this inefficiency underscores that overcoming the communication bottleneck is not merely an optimization, but a prerequisite for unlocking true scale of training across regions.
 Gradient compression and delayed aggregation are common strategies to improve the throughput of D-SGD in a lossy manner, but joint usage introduces a complex three-way trade-off.
-Gradient compression [^2] In this work, the gradient compression algorithm carries vanilla error-feedback (EF) stich2019error by default.
-EF is a mechanism that collects and reuses the errors from the gradient compression to retain the model accuracy.
-Other EF strategies, such as EF21 EF21 or EControl Stich2023EControl are out of this paper’s scope.
-reduces communication overhead by compressing gradient information before transmission, to mitigate the low bandwidth problem dgc .
-Delayed aggregation DGA applies updates using gradients computed several iterations earlier to alleviate the high latency problem.
-To overcome the problem of low-bandwidth and high-latency at the same time in WANs, we need to use these two strategies jointly, introducing intricate trade-offs in two key dimensions: (1) Iteration-to-accuracy: The impact of compressing delayed gradients on convergence remains uncertain, as it is unclear to what extent aggregating delayed compressed gradients will affect model accuracy, due to the lack of theoretical analysis in prior work.
+Gradient compression reduces communication overhead by compressing gradient information before transmission, to mitigate the low bandwidth problem yan2022acJsac.
+Delayed aggregation Chuanwu2022SAPipe; DGA applies updates using gradients computed several iterations earlier to alleviate the high latency problem.
+While combining these strategies offers a promising path to overcome the problem of low-bandwidth and high-latency at the same time, it introduces intricate trade-offs in two key dimensions: (1) Iteration-to-accuracy: The impact of compressing delayed gradients on convergence remains uncertain.
+It is unclear to what extent aggregating delayed compressed gradients will affect model accuracy, due to the lack of theoretical analysis in prior work.
 (2) Time-to-iteration: The effect of staleness and compression ratio on communication-computation overlap is not fully understood.
@@
 This threshold varies dynamically with the bandwidth and staleness.
-To achieve optimal time-to-accuracy, we analyze the above two dimensions separately.
+To minimize the wall-clock time required to reach the target accuracy, we analyze the above two dimensions separately.
 First, in iteration-to-accuracy, we propose a new analytical tool called the Nested Virtual Sequence.
@@
 With this tool, we are the first to derive the convergence rate of distributed SGD with delayed aggregation and gradient compression (DD-EF-SGD).
-This convergence rate establishes the iteration-to-accuracy relationship and reveals that the value of staleness exponentially worsens the negative effect of the gradient compression on the model convergence.
+This convergence rate establishes the iteration-to-accuracy relationship and reveals that increasing staleness exponentially worsens the negative effect of the gradient compression on the model convergence.
 Second, we use mathematical induction to model the relationship between the end-to-end time and the number of iterations in DD-EF-SGD.
-In this way, we build the time-to-accuracy relationship and propose an adaptive algorithm (shown in Fig.
-3) that can select the optimal compression ratio and staleness in the challenging dynamic network conditions, thus achieving the balanced three-way trade-off during the whole training.
-In a word, the contributions of our work are summarized as below:
+In this way, we build the time-to-accuracy relationship and propose an adaptive algorithm that can select the optimal compression ratio and staleness in the challenging dynamic network conditions, thus achieving the balanced three-way trade-off during the whole training.
+In summary, the contributions of our work are as follows:
   • We propose the Nested Virtual Sequence, a theoretical tool to analyze complex distributed SGD variants, by reducing it to a standard D-SGD process and several analytical noise terms.
-We are the first to derive the convergence rate of DD-EF-SGD with the tool, revealing that the delayed aggregation exponentially worsens the negative effect of the compression on the training process.
+We are the first to derive the convergence rate of DD-EF-SGD with the tool, revealing that delayed aggregation exponentially worsens the negative effect of the compression on the training process.
   • We mathematically establish an approximate relationship between end-to-end time and accuracy.
-Based on this, we propose DeCo-SGD, a joint optimization algorithm that adaptively calculates the optimal compression ratio and staleness according to the training environments.
-  • We apply DeCo-SGD  in challenging networks environments (average bandwidth < $1$Gbps, latency $\geq$ $100$ms) across diverse model architectures including CNN, GPT and ViT.
-Experimental results validate the great three-way trade-offs of our design, achieving up to 5.07×5.07$\times$5.07 × and 1.37×1.37$\times$1.37 × speed-ups over D-SGD and the SOTA static strategy, respectively.
+Based on this, we propose DeCo-SGD, a joint optimization algorithm that adaptively determines the optimal compression ratio and staleness under varying training environments.
+  • We apply DeCo-SGD  in challenging network environments (average bandwidth $\leq$ 11 Gbps, latency $\geq$ $100$ ms).
+We conduct extensive experiments that cover diverse model architectures (ResNet, GPT, ViT), various network conditions, and system scalability.
+Experimental results demonstrate that our design achieves up to $5.07\times$ and $1.37\times$ speed-ups over D-SGD and the SOTA static strategy, respectively.
 2 Preliminaries
-We specifically concentrate on distributed machine learning (DML) in data-parallel mode using the DD-EF-SGD as the optimizer.
-Figure 2: The running timelines for D-SGD and D-SGD variants.
-Both D-SGD and D-EF-SGD are serial processes, with D-EF-SGD achieving a reduced transmission time.
-Similarly, DD-SGD and DD-EF-SGD operate serially; while DD-SGD maintains the same communication time as D-SGD, DD-EF-SGD reduces transmission time.
+In this work, we focus on distributed machine learning (DML) in data-parallel mode using the DD-EF-SGD as the optimizer.
 2.1 The Optimization Problem of DML
 In the data-parallel paradigm, the training dataset is partitioned across $n$ worker nodes, each holding a subset of the data.
-The optimization problem of DML aims to minimize a global loss function $f(\cdot)$ based on the model parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $f(\cdot)$ is defined as a weighted sum of local loss functions of the $i$-th worker $f_{i}(\cdot)$:
+The optimization problem of DML aims to minimize a global loss function $f(\cdot)$ based on the model parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $f(\cdot)$ is defined as the average sum of local loss functions of the $i$-th worker $f_{i}(\cdot)$:
 $\underset{\mathbf{x}}{\min}\left[f(\mathbf{x}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})\right].$
@@
 2.2.1 D-SGD
-At $t$-th iteration, the $i$-th worker computes gradients $\mathbf{g}_{t}^{i}$ based on its local data and model parameters at $t$-th iteration $\mathbf{x}_{t}$.
+At the $t$-th iteration, the $i$-th worker computes gradients $\mathbf{g}_{t}^{i}$ based on its local data and model parameters at $t$-th iteration $\mathbf{x}_{t}$.
 $\gamma$ is the stepsize and these local gradients are aggregated to update the global model, following $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$.
 As shown in Fig.
-2, D-SGD is a serial computation process.
-In this process, computation and communication take place sequentially in each iteration, and communication includes end-to-end latency and transmission time.
-2.2.2 D-SGD with Gradient Compression and EF (D-EF-SGD)
-D-EF-SGD stich2020communication enhances D-SGD with the sparsification gradient compression [^3] In this work, the sparsification compressor in D-EF-SGD is Top-$k$ by default, a well-performed compressor in many hybrid gradient compressors stc ; CocktailSGD .
-by incorporating EF mechanisms to compensate for sparsification compression errors.
-Sparsification compression is $\mathbf{e}_{t}^{i}$ represents the error term of the $i$-th worker at the $t$-th iteration, storing the discrepancy between the compressed gradient and the true gradient.
-We denote $\mathbf{C}_{\delta}(\cdot)$ as the compression operator with the compression ratio $\delta$, $\Delta_{t}^{i}$ as the compressed update at $t$-th iteration in $i$-th worker, and the update rules for D-EF-SGD are as follows:
-$\begin{gathered}\Delta_{t}^{i}=\mathbf{C}_{\delta}(\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}),\quad\quad\quad\mathbf{e}_{t+1}^{i}=\mathbf{g}_{t}^{i}+\mathbf{e}_{t}^{i}-\Delta_{t}^{i},\quad\quad\quad\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\Delta_{t}^{i}.\end{gathered}$
-As shown in Fig. 2, D-EF-SGD retains a serial process, but reduces transmission time compared to D-SGD.
-2.2.3 D-SGD with Delay Aggregation (DD-SGD)
-DD-SGD DGA addresses communication delays by accounting for the delay staleness $\tau$ between the computation of gradients and their application to the model update stich2019error .
-$\mathbf{g}_{t-\tau}^{i}$ represents the gradient computed by worker $i$ at the $(t-\tau)$-th iteration and the update rule shows as $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}.$ The delay aggregation does not change the communication time for each iteration.
+3, standard D-SGD follows a strict sequential execution.
+The computation phase starts until the communication phase is fully completed, and communication includes end-to-end latency and transmission time.
+2.2.2 D-SGD with Gradient Compression
+Gradient compression is a typical communication optimization method to reduce the communication volumes during the training process by compressing transferred gradients cui2021slashing.
+Existing strategies generally fall into three categories: sparsification, quantization, and low-rank compression grace.
+Quantization compression lQuant2025jsac reduces the communication volume by decreasing the number of bits of gradients.
+However, it struggles to achieve aggressive compression ratios (such as $1\%$ or less) and lacks fine-grained adaptivity (like changing ratios from $12\%$ to $10\%$ or $4$-bit into $4.5$-bit quantization) due to its discrete value selection.
+Low-rank compression vogels2019powersgd is tailored for the all-reduce communication primitive and well suited for training within a single data center.
+However, it incurs higher compression costs compared to other compression strategies and cannot leverage its advantages in decentralized training, which prefers peer-to-peer communication under delayed aggregation strategy DGA.
+Sparsification compression transmits dagctmc only parts of gradients, offering two advantages for decentralized distributed training: (1) It enables aggressive compression of $1\%$, with the help of the error feedback (EF) mechanism [^2] In this work, the gradient compression algorithm carries vanilla EF stich2019error by default.
+EF is a mechanism that collects and reuses the errors from the gradient compression to retain the model accuracy.
+Other EF strategies, such as EF21 EF21 or EControl Stich2023EControl are out of this paper’s scope.; (2) Compression ratios are continuously selectable, facilitating easy adaptive adjustment.
+We use sparsification compression and incorporate EF.
+We denote $\mathbf{C}_{\delta}(\cdot)$ as the compression operator with the compression ratio $\delta$ and $\Delta_{t}^{i}$ as the compressed update at $t$-th iteration in $i$-th worker.
+$\mathbf{e}_{t}^{i}$ represents the error term of the $i$-th worker at the $t$-th iteration, storing the discrepancy between the compressed updates and the true ones.
+The update rules for D-SGD with gradient compression are $\Delta_{t}^{i}=\mathbf{C}_{\delta}(\mathbf{e}_{t}^{i}+\gamma\mathbf{g}_{t}^{i})$, $\mathbf{e}_{t+1}^{i}=\mathbf{e}_{t}^{i}+\gamma\mathbf{g}_{t}^{i}-\Delta_{t}^{i}$ and $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{1}{n}\sum_{i=1}^{n}\Delta_{t}^{i}$.
 As shown in Fig.
-2, DD-SGD parallelizes computation and communication by performing the $t$-th computation at the $t$-th iteration and sending $t$-th gradient, but the aggregation applies the gradient computed at $(t-\tau)$-th iteration.
-When combined with gradient compression (DD-EF-SGD), the transmission time is further reduced, as shown in the last part of Fig.
-2.
+3, D-SGD with gradient compression retains a serial process, but reduces transmission time compared to D-SGD.
+Figure 3: The running timelines for D-SGD and D-SGD with communication optimziation strategies.
+Both D-SGD and D-SGD with gradient compression are serial processes, with gradient compression reducing transmission time.
+D-SGD with delayed aggregation and DD-EF-SGD operate in parallel.
+2.2.3 D-SGD with Delayed Aggregation
+Delayed aggregation DGA addresses communication delays by accounting for the delay staleness $\tau$ between the computation of gradients and their application to the model update stich2019error.
+$\mathbf{g}_{t-\tau}^{i}$ represents the gradient computed by worker $i$ at the $(t-\tau)$-th iteration and the update rule shows as $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}$.
+Unlike standard D-SGD, delayed aggregation enables pipelining: the system performs the $t$-th computation concurrently with the transmission of gradients from the $(t-\tau)$-th iteration (as shown in Fig.
+3).
+While the physical transmission duration remains unchanged, it is effectively hidden behind the computation.
+When further combined with gradient compression (DD-EF-SGD), the transmission time is reduced.
 2.3 Virtual Sequence
-Virtual Sequence stich2018sparsified ; mania2017perturbed is an analytical tool that simplifies the convergence analysis of D-SGD variants like D-EF-SGD.
+The Virtual Sequence method stich2018sparsified; mania2017perturbed is an analytical tool designed to simplify the convergence analysis of D-SGD variants like D-SGD with gradient compression.
 It defines an auxiliary sequence, $\tilde{\mathbf{x}}_{t}$, which evolves according to an idealized update rule and serves as a reference for the actual iterates, $\mathbf{x}_{t}$, accounting for noise introduced by compression and delay.
@@
 Then we have $\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t+1}-B_{t+1}=\tilde{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\textbf{g}_{t}^{i}$.
-This can help theoretically analyze D-SGD variants, but fails in complex variants like DD-EF-SGD, because the noise term decoupled from DD-EF-SGD cannot be analyzed.
-A new method is needed to analyze these complex D-SGD variants.
+While effective for simple variants, this tool fails for complex coupled systems like DD-EF-SGD.
+The failure arises because the noise term in DD-EF-SGD involves a recursive coupling of compression errors and delay staleness.
+This dependency makes the noise term mathematically intractable under standard noise assumptions, necessitating a new analytical method to decouple these factors.
 3 Theoretical Analysis of DD-EF-SGD
-We derive the convergence rate of DD-EF-SGD. Notation list and detailed proof are in the Appendix.
+In this section, we establish the convergence guarantees for DD-EF-SGD.
+The key notations are summarized in Table 1.
+The training procedure of DD-EF-SGD is outlined in Algo.
+1 Detailed proofs are provided in the Appendix (Sec.
+B.2B.3).
+Table 1: Notation list.
+Notation Description
+$n$ number of workers
+$f_{i}(\cdot)$ the local loss function of worker $i$
+$f(\cdot)$ the global loss function, i.e., $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$
+$\mathbf{x}^{i}_{t}$
+the local model parameter in the $t$-th iteration
+of the $i$-th worker
+$\mathbf{x}_{t}$ the global model parameter in the $t$-th iteration
+$\mathbf{g}^{i}_{t}$ the stochastic gradient of $\textbf{x}_{t}$ of the worker $i$
+$\mathbf{g}_{t}$ the average of $\mathbf{g}^{i}_{t}$, i.e., $\mathbf{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}^{i}_{t}$
+$\gamma$ the stepsize
+$\mathbf{T}$ total number of iterations
+$\delta$ compression ratio ($0<\delta\leq 1$)
+$\mathbf{C}_{\delta}(\cdot)$
+the sparsification compressor with
+compression ratio $\delta$
+$\tau$ the delayed staleness
+$\mathbf{e}_{t}^{i}$
+the local error term of the $i$-th worker
+in the $t$-th iteration
+$V_{g}$ the volume of the gradient (bits)
+$E$ the frequency to update $\tau$ and $\delta$ with DeCo
+$\bar{a}$
+the average bandwidth (bits/s) of the slowest link
+in previous $E$ iterations
+$b$ the network end-to-end latency (s) of the slowest link
+$T_{\text{comp}}$ computation time per iteration (s)
+$T_{\text{avg}}$ average end-to-end training time per iteration (s)
+$TS_{t}$ end time of the $t$-th computation (s)
+$TC_{t}$ end time of the $t$-th communication (s)
+Input: $n$, $\gamma$, $\mathbf{T}$, $C(\cdot)$, $\delta$, $\tau$
+Output: $\textbf{x}_{\mathbf{T}}$
+Initialize $\mathbf{x}_{0}$, $\mathbf{e}_{0}^{i}=\mathbf{0}_{d}$;
+for $t\in[\mathbf{T}]$ do
+ /* Worker side */
+ for $i\in[n]$ do
+      Worker $i$ waits until it receives $\textbf{x}_{t}$;
+      Calculate $\mathbf{g}_{t}^{i}$ based on the global model $\mathbf{x}_{t}$;
+    $\Delta_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i})$;
+    $\textbf{e}_{t+1}^{i}=\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i}-\Delta_{t}^{i}$;
+      Send $\Delta_{t}^{i}$ to the server (non-blocking);
+   end for
+ /* Server side */
+   Server waits until it receives $\Delta_{t-\tau}^{i}$ for all $i\in[n]$;
+ $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{1}{n}\sum_{i=1}^{n}\Delta_{t-\tau}^{i}$;
+   Broadcast $\textbf{x}_{t+1}$;
+end for
+Return $\textbf{x}_{\mathbf{T}}$;
+Algorithm 1 $D$-EF-SGD
 3.1 Regular Assumptions
-We assume functions are $L$-smooth with the gradient noise of D-SGD presumed to exhibit zero mean and a variance of $\sigma^{2}$.
-We gauge data heterogeneity using constants $\zeta_{i}^{2}>0$, which cap the variance among the $n$ workers.
-In Theorem 2, we posit that objective functions are $\mu$-strongly convex.
-Detailed assumptions are shown below.
+We adopt standard assumptions, following previous works cui2022infocom; stich2020communication.
 Assumption 1 ($L$-smoothness). We assume $L$-smoothness of $f_{i},\ i\in[n]$, that is for all $\mathbf{x},\ \textbf{y}\in\mathbb{R}^{d}$:
 $\lVert\nabla f_{i}(\textbf{y})-\nabla f_{i}(\mathbf{x})\rVert\leq L\lVert\textbf{y}-\mathbf{x}\rVert.$ (1)
-Assumption 2 ($\mu$-strongly convexity). We assume $\mu$-strong convexity of $f_{i},i\in[n]$, that is for all $\mathbf{x},\textbf{y}\in\mathbb{R}^{d}$:
-$f_{i}(\mathbf{x})-f_{i}(\textbf{y})\geq\langle\nabla f_{i}(\textbf{y}),\mathbf{x}-\textbf{y}\rangle+\frac{\mu}{2}\lVert\nabla f_{i}(\mathbf{x})-\nabla f_{i}(\textbf{y})\rVert^{2}.$ (2)
 Assumption 2 (Bounded gradient noise).
@@
 We define the stochastic gradient noise of $i$-th worker as $\xi^{i}$ and for simplicity, we only consider the instructive case of uniformly bounded $\xi^{i}$ for all $\mathbf{x}\in\mathbb{R}^{d},\ i\in[n]$:
-$\mathbf{g}^{i}_{t}=\nabla f_{i}(\mathbf{x}_{t})+\boldsymbol{\xi}^{i},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\boldsymbol{\xi}^{i}=\textbf{0}_{d},\quad\quad\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\lVert\boldsymbol{\xi}^{i}\rVert^{2}\leq\sigma^{2}.$ (3)
+$\mathbf{g}^{i}_{t}=\nabla f_{i}(\mathbf{x}_{t})+\boldsymbol{\xi}^{i},\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\boldsymbol{\xi}^{i}=\textbf{0}_{d},\quad\mathbb{E}_{\boldsymbol{\xi}^{i}}\lVert\boldsymbol{\xi}^{i}\rVert^{2}\leq\sigma^{2}.$ (2)
 Assumption 3 (Measurement of data heterogeneity). We measure data dissimilarity by constants $\zeta_{i}^{2}\geq 0$ that bound the variance across the $n$ nodes. We have:
-$\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+\lVert\nabla f(\mathbf{x})\rVert^{2},\quad\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2},\quad\forall\mathbf{x}\in\mathbb{R}^{d},i\in[n].$ (4)
-These assumptions are standard, following previous works cui2022infocom ; stich2020communication .
-For Assumption 3, we can assume that the $\sigma^{2}$ increases with the gradient norm hardthreshold .
-In this way, by replacing $\sigma$ with a new $\sigma^{\prime}$, one can obtain a tighter bound.
-Similarly in Assumption 4, the general version is $\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta_{i}^{2}+Z\lVert\nabla f(\mathbf{x})\rVert^{2}$, and we take $Z=1$ by default.
-One can take a more general bound by adding $Z$.
-However, extending the analysis to cover these generalized assumptions is beyond the scope of this work.
-3.2 Nested Virtual Sequences (NVS)
-To solve the problem that the traditional Virtual Sequence method cannot analyze complex D-SGD variants like DD-EF-SGD. We propose NVS, which is composed of two parts. In the first part, we have
+$\frac{1}{n}\sum_{i\in[n]}\lVert\nabla f_{i}(\mathbf{x})\rVert^{2}\leq\zeta^{2}+Z^{2}\lVert\nabla f(\mathbf{x})\rVert^{2},\forall\mathbf{x}\in\mathbb{R}^{d}.$ (3)
+Under strongly convex objective functions, we use Assumption 4 in addition.
+Assumption 4 ($\mu$-strongly convex). We assume $\mu$-strong convexity of $f_{i},i\in[n]$, that is for all $\mathbf{x},\textbf{y}\in\mathbb{R}^{d}$:
+$f_{i}(\mathbf{x})-f_{i}(\textbf{y})\geq\langle\nabla f_{i}(\textbf{y}),\mathbf{x}-\textbf{y}\rangle+\frac{\mu}{2}\lVert\nabla f_{i}(\mathbf{x})-\nabla f_{i}(\textbf{y})\rVert^{2}.$ (4)
+3.2 Theoretical Tool: Nested Virtual Sequences (NVS)
+A major challenge in analyzing DD-EF-SGD is the coupling effect between compression error and delay staleness.
+Traditional Virtual Sequence methods fail here because the noise term becomes recursively dependent on historical states, making it mathematically intractable.
+We propose NVS, which is composed of two parts. In the first part, we have
 $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\mathbf{v}_{t},\quad\quad\quad B_{t+1}=B_{t}+\tilde{\mathbf{v}}_{t}-\mathbf{v}_{t}.$ (5)
@@
 In the second part, we use the virtual sequence framework based on the new items $\tilde{\mathbf{x}}_{t}$ and $\tilde{\mathbf{v}}_{t}$.
-We define the second noise term as $\tilde{B}_{t}$ and have:
-$\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\tilde{B}_{t+1}=\tilde{B}_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\tilde{\mathbf{v}}_{t},\quad\quad\quad\hat{\mathbf{x}}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{B}_{t}.$
+We define the second noise term as $\tilde{B}_{t}$, having $\tilde{\mathbf{x}}_{t+1}=\tilde{\mathbf{x}}_{t}-\tilde{\mathbf{v}}_{t}$, $\tilde{B}_{t+1}=\tilde{B}_{t}+\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}-\tilde{\mathbf{v}}_{t}$ and $\hat{\mathbf{x}}_{t}=\tilde{\mathbf{x}}_{t}-\tilde{B}_{t}$.
 Similarly, it is easy to get $\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$.
 In this way, we decouple DD-EF-SGD into a standard D-SGD process and two analyzable noise terms:
-$\mathbf{v}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{C}_{\delta}(\mathbf{g}_{t-\tau}^{i}+e_{t-\tau}^{i}),\quad B_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}e_{t-\tau}^{i},$
-$\tilde{\mathbf{v}}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i},\quad\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-j}^{i}.$ (6)
+$\mathbf{v}_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{C}_{\delta}(\gamma\mathbf{g}_{t-\tau}^{i}+\mathbf{e}_{t-\tau}^{i}),$ $B_{t}$ $=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{e}_{t-\tau}^{i},$ (6)
+$\tilde{\mathbf{v}}_{t}$ $=\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i},$ $\tilde{B}_{t}$ $=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-j}^{i}.$ (7)
+For form consistency in all iterative expressions, we define $\mathbf{g}_{t-\tau}^{i}=\mathbf{e}_{t-\tau}^{i}=\mathbf{0}^{d}$, for all $i\in[n]$, if $t-\tau\leq 0$.
+This means that when $t\leq\tau$, we have
+$\mathbf{v}_{t}=\tilde{\mathbf{v}}_{t}=B_{t}=\mathbf{0}^{d},\quad\quad\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{t}\mathbf{g}_{t-j}^{i}.$
+It is worth noting that since $\tau$ is typically much smaller than the total number of training iterations $T$, the case $t\leq\tau$ only occurs at the early stage of training.
+Consequently, it has a negligible influence on the overall convergence analysis and we will not discuss this boundary case.
 3.3 Convergence Rate of DD-EF-SGD
+Table 2: Summary of the convergence results for different algorithms under non-convex and $\mu$-strongly convex cases.
+We define $\phi=\frac{1-\delta}{\delta\,(1-\tfrac{\delta}{2})^{\tau}}$, $F_{0}\geq f(\mathbf{x}_{0})-f^{*}$.
+Algorithm Non-convex $\mu$-strongly convex
+$D$-EF-SGD $\begin{aligned} \mathcal{O}(&\frac{\sigma^{2}}{n\,\epsilon^{2}}+\frac{\sqrt{\;\tfrac{\phi\,\zeta^{2}}{\delta}+\left(\phi+\tfrac{\tau}{n}\right)\sigma^{2}\;}}{\epsilon^{3/2}}+\\[-3.0pt] &\frac{\tau+Z\sqrt{\phi/\delta}}{\epsilon})\cdot LF_{0}\end{aligned}$ $\begin{aligned} \mathcal{O}(&\frac{\sigma^{2}}{n\,\mu\,\epsilon}+\frac{\sqrt{\,L\!\left(\tfrac{\phi\,\zeta^{2}}{\delta}+\left(\phi+\tfrac{\tau}{n}\right)\sigma^{2}\right)\,}}{\mu\,\epsilon^{1/2}}+\\[-3.0pt] &\frac{L(\tau+Z\sqrt{\phi/\delta})}{\mu})\end{aligned}$
+D-SGD with delayed aggregation DGA $\mathcal{O}(\frac{\sigma^{2}}{n\,\epsilon^{2}}+\frac{\sqrt{\;\tfrac{\tau}{n}\,\sigma^{2}\;}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon})\cdot LF_{0}$ $\mathcal{O}(\frac{\sigma^{2}}{n\,\mu\,\epsilon}+\frac{\sqrt{\,L\,\tfrac{\tau}{n}\,\sigma^{2}\,}}{\mu\,\epsilon^{1/2}}+\frac{L\tau}{\mu})$
+D-SGD with gradient compression stich2020communication $\mathcal{O}(\frac{\sigma^{2}}{n\,\epsilon^{2}}+\frac{\sqrt{\;\tfrac{\zeta^{2}}{\delta}+\sigma^{2}\;}}{\sqrt{\delta}\,\epsilon^{3/2}}+\frac{Z}{\delta\,\epsilon})\cdot LF_{0}$ $\mathcal{O}(\frac{\sigma^{2}}{n\,\mu\,\epsilon}+\frac{\sqrt{\,L\!\left(\tfrac{\zeta^{2}}{\delta}+\sigma^{2}\right)\,}}{\sqrt{\delta}\,\mu\,\epsilon^{1/2}}+\frac{LZ}{\mu\,\delta})$
 Theorem 1 (Non-convex convergence rate of DD-EF-SGD).
 Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth.
-There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
-$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{1}{\epsilon}+\frac{\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$ (7)
+There exists a stepsize $\gamma\leq\min\{\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
+$\begin{split}\mathcal{O}\Bigg(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\tfrac{\phi\zeta^{2}}{\delta}+\left(\phi+\tfrac{\tau}{n}\right)\sigma^{2}}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon}+\frac{Z\sqrt{\phi}}{\sqrt{\delta}\,\epsilon}\Bigg)\cdot L\,(f(\mathbf{x}_{0})-f^{*})\end{split}$
 iterations of DD-EF-SGD, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total number of iterations, chosen at random uniformly.
-Remark 1.
-($\phi$ determining the convergence in the non-degradation condition) In typical DML scenarios, $\frac{\tau}{n}\leq 1$.
+Remark 1 ($\phi$ determining the convergence in the non-degradation condition).
+In typical distributed training scenarios, the number of workers $n$ is large and $\frac{\tau}{n}\leq 1$.
 As the training scale increases, $\tau$ tends to decrease due to the larger model requiring larger computation time, and $n$ tends to increase, resulting in an even smaller $\frac{\tau}{n}$.
-While to reduce the communication traffic, $\delta$ tends to be $10\%$ or less, common in DML CocktailSGD ; grace , leading to the order of magnitude of $\phi$ larger than that of $\frac{\tau}{n}$.
-Then the convergence rate is $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}+\frac{\sqrt{\frac{\phi}{\delta}}+\sqrt{\tau}}{\epsilon})$.
-Similar to prior works dagctmc , the frist term $\frac{\sigma^{2}}{n\epsilon^{2}}$ is independent of $\delta$ and $\tau$.
+While to reduce the communication traffic, $\delta$ tends to be $10\%$ or less, common in training with gradient compression dgc; grace, leading to the order of magnitude of $\phi$ larger than that of $\frac{\tau}{n}$.
+Then the convergence rate can be written as $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}+\frac{Z\sqrt{\frac{\phi}{\delta}}+\tau}{\epsilon})$.
+Similarly to previous work dagctmc, the frist term $\frac{\sigma^{2}}{n\epsilon^{2}}$ is independent of $\delta$ and $\tau$.
 The second term $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$ has a greater impact on the overall convergence rate compared to the third term.
-Let us focus on $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$, in LLMs pre-training scenarios, local datasets are typically allocated centrally, resulting in low data heterogeneity $\zeta$.
+Let us focus on $\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\epsilon^{3/2}}$, in LLMs training scenarios (like pre-training or full-parameter fine-tuning), local datasets are typically allocated centrally, resulting in IID datasets ($\zeta\approx 0$ ).
 Moreover, due to the small batch sizes commonly used in LLM training (limited by the constraint of GPU memory), $\sigma$ tends to be large.
 Under such circumstances, $\frac{\zeta^{2}}{\delta}\ll\sigma^{2}$ and $\phi$ captures the convergence rate.
-While in Federated Learning (large $\zeta$) or training small models (small $\sigma^{2}$), the key factor is $\phi^{\prime}=\frac{1-\delta}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}$.
+While in non-IID scenarios, (large $\zeta$) or training small models (small $\sigma^{2}$), the key factor is $\phi^{\prime}=\frac{1-\delta}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}$.
 In this work, we focus on the first scenario where $\phi$ is the key factor that determines the convergence of DD-EF-SGD.
-For the analysis in Federated Learning or training small models, simply replacing $\phi$ with $\phi^{\prime}$ is enough and we do not discuss this later in the paper.
-Remark 2.
-(Degradation condition) DD-EF-SGD has two degradation conditions, namely, without compression ($\delta=1$) and without delayed aggregation ($\tau=0$).
-When $\delta=1$, $\phi=0$ and DD-EF-SGD degrades into DD-SGD.
-Then the convergence rate is equal to $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\sqrt{\tau}}{\epsilon})$, same with the convergence rate of DD-SGD stich2022sharper ; stich2019error .
-When $\tau=0$, DD-EF-SGD degrades to D-EF-SGD, and $\phi=\frac{1-\delta}{\delta}$.
-Then the convergence rate becomes $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\zeta/\sqrt{\delta}+\sigma}{\sqrt{\delta}\epsilon^{3/2}}+\frac{L}{\delta\epsilon})$, same with the convergence rate of D-EF-SGD stich2020communication ; dagc .
+For the analysis in non-IID scenarios (like Federated Learning), we use $\phi^{\prime}$ as the key factor.
+Remark 2 (Degradation condition).
+DD-EF-SGD has two degradation conditions, namely, without compression ($\delta=1$) and without delayed aggregation ($\tau=0$).
+When $\delta=1$, $\phi=0$ and DD-EF-SGD degrades into D-SGD with delayed aggregation.
+Then the convergence rate is equal to $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon})$, same with the convergence rate of D-SGD with delayed aggregation stich2022sharper; stich2019error.
+When $\tau=0$, DD-EF-SGD degrades to D-SGD with gradient compression, and $\phi=\frac{1-\delta}{\delta}$.
+Then the convergence rate becomes $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\zeta/\sqrt{\delta}+\sigma}{\sqrt{\delta}\epsilon^{3/2}}+\frac{Z}{\delta\epsilon})$, same with the convergence rate of D-SGD with gradient compression stich2020communication.
 Theorem 2 (Convex convergence rate of DD-EF-SGD).
 Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex.
-Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
-$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$ (8)
-iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
+Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4\sqrt{2}L\tau},\frac{1}{16LZ\sqrt{2\phi/\delta}}\}$, such that at most
+$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L\tau}{\mu}+\frac{LZ\sqrt{\phi/\delta}}{\mu})\end{split}$
+iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, selected probabilistically based on $(1-\frac{\mu\gamma}{2})^{-t}$.
 Remark 3.
-Same to remarks in Theorem 1, in the degradation conditions, the convex convergence rate is the same as that of DD-SGD (under $\delta=1$) and D-EF-SGD (under $\tau=0$).
-In the non-degradation conditions, the convergence rate can be written into $\mathcal{O}(\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{\sqrt{\frac{L\phi}{\delta}}+\sqrt{L\tau}}{\mu})$, $\phi$ dominates the convergence rate of DD-EF-SGD in the convex case, too.
+Same to remarks in Theorem 1, in the degradation conditions, the convex convergence rate is the same as that of D-SGD with delayed aggregation (under $\delta=1$) and D-SGD with gradient compression (under $\tau=0$).
+In the non-degradation conditions, the convergence rate can be written into $\mathcal{O}(\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{\phi(\frac{\zeta^{2}}{\delta}+\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L(\tau+Z\sqrt{\phi/\delta})}{\mu})$, $\phi$ dominates the convergence rate of DD-EF-SGD in the convex case, too.
 4 Mathematical Modeling Time to Accuracy in DD-EF-SGD
-[…]
+We denote the bandwidth (bits/s) as a time-varying function $a(s)$, where $s$ represents continuous time, the average bandwidth over the previous $E$ iterations as $\bar{a}$, end-to-end latency (s) as $b$, and gradient size (bits) as $V_{g}$.
+We denote the computation time per iteration (s) and the average end-to-end training time per iteration (s) as $T_{\text{comp}}$ and $T_{\text{avg}}$, respectively.
+We use $\lceil\cdot\rceil$ to represent the ceiling function.
+Detailed proofs are provided in the Appendix.
+4.1 Iteration Timing Model of DD-EF-SGD
+In DD-EF-SGD with one worker and server, each iteration consists of three stages: local computation $T_{\text{comp}}$; gradient transmission time $t_{\delta}$, where $\int_{0}^{t_{\delta}}a(s)\,\mathrm{d}s{=}\delta V_{g}$ and fixed end-to-end latency $b$.
+Let $t$ denote the iteration index ($t=0,1,2,\dots$).
+Considering the timing constraint under delayed aggregation DGA, the (t+1)(t{+}1)-th computation starts after both the $t$-th computation and the $(t{-}\tau)$-th update are completed.
+Let $TS_{t}$ and $TC_{t}$ be the completion times of the $t$-th computation and communication, respectively.
+This leads to the following recursive relation:
+$TS_{t+1}=T_{\text{comp}}+\max\{\,TS_{t},\;TC_{t-\tau}\,\}.$ (8)
+4.2 Modeling Assumptions
+To facilitate the time-to-iteration model in DD-EF-SGD  we introduce the following assumptions.
+Assumption 5 (Bottleneck equivalence).
+In synchronous parameter-server systems, the overall iteration time is dominated by the slowest link.
+This allows us to analyze the multi-worker system by focusing on a single worker-server link representing the bottleneck, without loss of generality for the analysis.
+Assumption 6 (Historical average bandwidth estimation).
+Based on network measurement studies strauss2003measurement; li2004first, network bandwidth exhibits local stationarity at minute-level timescales, with average bandwidth prediction errors typically within $10\%$ ng2002predicting.
+We define the window for recording bandwidth as $E$ iterations.
+Therefore, we let $a(s)=\bar{a}$ in the following $E$ iterations, Every $E$ iterations, we update the bandwidth.
+That means the gradient transmission time $t_{\delta}{=}\frac{\delta V_{g}}{\bar{a}}$.
 4.3 Time to Iteration in DD-EF-SGD
-Theorem 3 (Estimation of $T_{avg}$ in DD-EF-SGD).
-Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed.
-Then the average iteration time $T_{avg}$ satisfies the following approximation:
-$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\delta\cdot S_{g}/a}{\tau+1},\frac{\delta\cdot S_{g}}{a},T_{\text{comp}}\right\}},$ (9)
-where the approximation error is upper bounded by $O(\frac{1}{t})$.
-Remark 4 (The optimal compression ratio $\delta^{*}$).
-We can minimize $T_{avg}$ to $T_{comp}$ by enlarging $\tau$ and choosing a small $\delta$ Let $\delta^{*}(\tau)=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$, we can obtain that $\delta\leq\delta^{*}(\tau)$.
-Given $\tau\geq\frac{b}{T_{comp}}$, $\delta^{*}=\delta^{*}(\tau)$ This is due to the fact that the more aggressive $\delta$ cannot get the performance gain, which is covered by the delay while leading to accuracy degradation.
-In this way, $\delta^{*}$ is the maximal value in the given range.
+Theorem 3 (Estimation of $T_{\text{avg}}$ in DD-EF-SGD).
+Under Assumptions 5, 6, the average iteration time $T_{\text{avg}}$ satisfies the following approximation:
+$T_{\text{avg}}\approx\max{\left\{\frac{T_{\text{comp}}+b+\delta\cdot V_{g}/\bar{a}}{\tau+1},\frac{\delta\cdot V_{g}}{\bar{a}},T_{\text{comp}}\right\}},$ (9)
+where the approximation error is caused by the pipeline boundary and can be neglected when $E\gg\tau$.
+Remark 4 (The locally optimal compression ratio $\delta^{*}(\tau)$).
+Our goal is to minimize the iteration time $T_{\text{avg}}$ without strictly sacrificing model accuracy.
+According to Eq.
+(9), the lower bound of the average iteration time is the computation time $T_{\text{comp}}$.
+There exists a locally optimal compression ratio given a fixed $\tau$, denoted as $\delta^{*}(\tau)$, such that $T_{\text{avg}}(\tau,\delta^{*}(\tau))=T_{\text{comp}}$.
+Choosing any $\delta\leq\delta^{*}$ yields no gain in $T_{\text{avg}}$, but degrade the model performance due to more information loss.
+Thus, for a $\tau$, $\delta^{*}(\tau)=\min\left\{\max\left(0,(\tau T_{\text{comp}}-b)\frac{\bar{a}}{V_{g}}\right),T_{\text{comp}}\frac{\bar{a}}{V_{g}},1\right\}$.
 4.4 DeCo: Joint Optimization for Delay Staleness and Gradient Compression in DML
@@
 DeCo-SGD adaptively adjusts the compression ratio $\delta$ and delay staleness $\tau$ based on the dynamic network conditions monitored by the worker.
-Input: $S_{g}$, $a$, $b$, $T_{comp}$
+Input: $n$, $\gamma$, $\mathbf{T}$, $C(\cdot)$, $V_{g}$, $T_{\text{comp}}$, $E$
+Output: $\textbf{x}_{\mathbf{T}}$
+Initialize $\mathbf{x}_{0}$, $\mathbf{e}_{0}^{i}=\mathbf{0}_{d}$;
+Initialize $\tau$ and $\delta$ based on DeCo algorithm;
+for $t\in[\mathbf{T}]$ do
+ /* Worker side */
+ for $i\in[n]$ do
+      Worker $i$ waits until it receives $\textbf{x}_{t}$;
+      Calculate $\mathbf{g}_{t}^{i}$ based on the global model $\mathbf{x}_{t}$;
+    $\Delta_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i})$;
+    $\textbf{e}_{t+1}^{i}=\textbf{e}_{t}^{i}+\gamma\textbf{g}_{t}^{i}-\Delta_{t}^{i}$;
+      Send $\Delta_{t}^{i}$ to the server;
+   end for
+ /* Server side */
+   Server waits until it receives $\Delta_{t-\tau}^{i}$ for all $i\in[n]$;
+ $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{1}{n}\sum_{i=1}^{n}\Delta_{t-\tau}^{i}$;
+   Broadcast $\textbf{x}_{t+1}$;
+ /* Update $\tau$ and $\delta$ with DeCo */
+ if $(t+1)\mod E==0$ then
+      Get $\bar{a}$, $b$ from the network;
+    $\tau$, $\delta$ = DeCo($V_{g}$, $\bar{a}$, $b$, $T_{\text{comp}}$);
+      Broadcast new $\tau$ and $\delta$ to workers;
+   end if
+end for
+Return $\textbf{x}_{\mathbf{T}}$;
+Algorithm 2 DeCo-SGD
+Input: $V_{g}$, $\bar{a}$, $b$, $T_{\text{comp}}$
 Output: Optimal $\tau$, $\delta$ with minimal $\phi$
 $\phi_{min}=\infty$
-for $\tau\in\{\lceil\frac{b+S_{g}/a}{T_{comp}}\rceil,\cdots,1+\lceil\frac{b}{T_{comp}}\rceil,\lceil\frac{b}{T_{comp}}\rceil\}$  do
-      $\delta=\min\left\{(\tau T_{comp}-b)\frac{a}{S_{g}},T_{comp}\frac{a}{S_{g}},1\right\}$;
-     /* Calculate the value of $\phi$/
+for $\tau\in\{\lceil\frac{b+V_{g}/\bar{a}}{T_{\text{comp}}}\rceil,\cdots,1+\lceil\frac{b}{T_{\text{comp}}}\rceil,\lceil\frac{b}{T_{\text{comp}}}\rceil\}$ do
+   $\delta=\min\left\{(\tau T_{\text{comp}}-b)\frac{\bar{a}}{V_{g}},T_{\text{comp}}\frac{\bar{a}}{V_{g}},1\right\}$;
+  /* Calculate the value of $\phi$ */
    $\phi=\frac{1-\delta}{\delta\cdot(1-\frac{\delta}{2})^{\tau}}$;
@@
 Algorithm 3 DeCo
-Input: $n$, $\gamma$, $S_{g}$, $T_{comp}$, initial parameters $\textbf{x}_{0}$, initial local error $\textbf{e}_{0}^{i}=\textbf{0}_{d}$, the frequency to update DeCo $E$
-Output: $\textbf{x}_{T}$
-for $i=1,\ldots,n$ do
-       if $t\mod E==1$ then
-             /* Update $\tau$ and $\delta$ with DeCo per $E$ iterations */
-             Get $a$, $b$ from the network;
-             $\tau$, $\delta$ = DeCo($S_{g}$, $a$, $b$, $T_{comp}$);
-       end if
-      $\hat{\Delta}_{t}^{i}=\textbf{C}_{\delta}(\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i})$;
-       $\textbf{e}_{t+1-\tau}^{i}=\textbf{e}_{t-\tau}^{i}+\textbf{g}_{t-\tau}^{i}-\hat{\Delta}_{t}^{i}$;
-       $\textbf{x}_{t+1}=\textbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\hat{\Delta}_{t}^{i}$;
-end for
-Return $\textbf{x}_{T}$;
-Algorithm 2 DeCo-SGD
 Our design goal is to get the minimal end-to-end time to achieve a specific accuracy, which is equivalent to solving the minimum of the binary function $\phi$.
 Since $\tau$ is discrete, the minimum of $\phi$ cannot be directly solved by the partial derivatives.
-We use the critical condition $T_{avg}=T_{comp}$ (i.e., there are no redundant bubbles in the pipeline) to reduce the global optima into a countable number of local optima.
+We use the critical condition $T_{\text{avg}}=T_{comp}$ (i.e., there are no redundant bubbles in the pipeline) to reduce the global optima into a countable number of local optima.
 The mathematical description of this goal is
-$\min\ \phi(\delta,\tau)=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}},\quad\textit{s.t.}\ T_{avg}=T_{comp},\quad\tau\in\mathbb{N},\quad\delta\in\left(0,1\right].$ (10)
+$\min\ \phi(\delta,\tau)=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}},\textit{s.t.}\ T_{\text{avg}}{=}T_{\text{comp}},\tau\in\mathbb{N},\delta\in\left(0,1\right].$ (10)
 Combined with Remark 4, we can rewrite Eq. 10 into:
-$\min\ \phi(\delta^{*}(\tau),\tau)=\frac{1-\delta^{*}(\tau)}{\delta^{*}(\tau)(1-\frac{\delta^{*}(\tau)}{2})^{\tau}}\quad\textit{s.t.}\ \tau\in\left[\lceil\frac{b}{T_{comp}}\rceil,\cdots,\lceil\frac{b+\frac{S_{g}}{a}}{T_{comp}}\rceil\right].$ (11)
-We traverse the range of $\tau$ to get $\phi^{*}$ to achieve the fastest convergence rate and choose the smallest $\tau$ as $\tau^{*}$ The procedure for obtaining $\tau^{*}$ and $\delta^{*}$ is referred to as the DeCo algorithm (shown in Algo.
-1).
-The implementation of DeCo in D-SGD is called DeCo-SGD (pseudo-code shown in Algo.
-2 and design shown in Fig.
+$\min\ \phi(\delta^{*}(\tau),\tau)=\frac{1-\delta^{*}(\tau)}{\delta^{*}(\tau)(1-\frac{\delta^{*}(\tau)}{2})^{\tau}}$ (11)
+$\textit{s.t.}\ \tau\in\left[\lceil\frac{b}{T_{comp}}\rceil,\cdots,\lceil\frac{b+\frac{V_{g}}{\bar{a}}}{T_{comp}}\rceil\right].$
+We traverse the range of $\tau$ to get $\phi^{*}$ to achieve the fastest convergence rate and choose the smallest $\tau$ as $\tau^{*}$ .
+The procedure for obtaining $\tau^{*}$ and $\delta^{*}$ is referred to as the DeCo algorithm (shown in Algo.
 3).
-DeCo-SGD sets a frequency $E$ to update DeCo.
-$E$ describes the sensitivity of DeCo.
-To enable frequent adjustments of the compression ratio and delay staleness in response to a dynamic network environment, a smaller $E$ value can be adopted, such as $E=1$.
-In the normal case, a middle $E$ like $100$ is enough.
-DeCo-SGD incurs a negligible extra cost compared to vanilla DD-EF-SGD and the extra time complexity is $\mathcal{O}(\frac{T}{E})$.
-The increase in $n$ does not result in additional cost.
+Figure 5: System implementation of DeCo-SGD.
+Each worker has: (1) Local compute unit to local compute and compress gradients; (2) Network monitor to get the bandwidth and latency; The server has a cache to save staled updates, the compute unit to update the model based on staled updates and DeCo Unit to decide ($\delta,\tau$) based on the network conditions.
+Workers and the server communicate with each other with Comm.
+API.
+4.5 Efficient Implementation of DeCo
+The implementation of DeCo in D-SGD is integrated into DeCo-SGD (pseudo-code shown in Algo.
+2, design in Fig.
+4, and system implementation in Fig.
+5).
+We use a pre-computed Look-up Table (LUT) strategy to ensure its decision-making process runs in $\mathcal{O}(1)$ runtime complexity.
+The implementation consists of two operational phases:
+4.5.1 Pre-computation
+To build the LUT, DeCo-SGD profiles the network during a warm-up phase to estimate the bandwidth range $a_{min},a_{max}]$.
+This range is then linearly interpolated into several sample points (like $20$ points).
+With measured $b$, the server runs DeCo offline for each sample point to compute and store the optimal configuration.
+4.5.2 Runtime Lookup
+During training, the worker updates the average bandwidth $\bar{a}$ per $E$ iterations.
+The system then queries the LUT for the optimal configuration.
+If $\bar{a}$ is within the recorded range, the system retrieves $(\delta^{*},\tau^{*})$ by rounding $\bar{a}$ to the nearest discretized bandwidth in the LUT.
+This $O(1)$ lookup ensures minimal runtime cost.
+If $\bar{a}$ falls outside the recorded range $a_{\min},a_{\max}]$, DeCo-SGD falls back to online computation in the server to ensure stability and correctness.
 5 Evaluation Experiments
+5.1 Experimental Environment
+The experiments are conducted on a distributed cluster comprising $32$ nodes.
+Each node is equipped with an Ubuntu 24.04 LTS system, an Intel Xeon Gold $6230$ CPU, and a single Nvidia A4040 GPU with 4848GB memory.
+The Python version is $3.10.16$, and other used libraries are all based on the Python version.
+We use Python $3.10.16$, PyTorch $2.5.1$ with CUDA $12.4$ as the ML toolkit and Gloo as the communication backend.
 5.2 Experimental Settings
-We evaluate our approach on diverse CV and NLP tasks using CNN lecun1998gradient , ViT vit , and GPT gpt2 models, including FashionMNIST xiao2017fashion , CIFAR-$10$, ImageNet imagenet , and Wikitext wikitext .
-Our baselines include D-SGD, Accordion accordion , DGA DGA , and CocktailSGD CocktailSGD .
-CocktailSGD is a SOTA static strategy that combines gradient compression and delayed aggregation using fixed parameters.
-If not specified, all compressors are Top-$k$ gradient compression dagctmc , except for CocktailSGD.
+Experiment tasks: To demonstrate the superiority of our design, we evaluate three representative tasks: CIFAR-$10$ with ResNet-34 resnet (denoted as ResNet@CIFAR-$10$), ImageNet imagenet using ViT vit (denoted as ViT@ImageNet) and Wikitext wikitext on GPT gpt2 (denoted as GPT@Wikitext).
+The ResNet architecture is ResNet-34, used widely in previous works dagctmc; stc.
+The ViT base model features approximately $86$ million parameters, while the smallest GPT variant contains $124$ million parameters.
+For ResNet@CIFAR-$10$, the learning rate is $0.01$, and the batch size per node is $32$.
+For ViT@ImageNet and GPT@Wikitext, the learning rates are $0.01$ and $0.1$, respectively, with batch sizes of $40$ and $5$.
+Regarding the initialization strategy, we train ResNet from scratch, whereas ViT and GPT are initialized with pre-trained weights.
+Baselines: We compare D-SGD, Accordion, DAGC, DC2, DGA, CocktailSGD as baselines.
+D-SGD is the case $\tau=0,\delta=1$ .
+DAGC dagctmc , Accordion accordion, DC2 dc2 are related adaptive gradient compression algorithms.
+DAGC is the SOTA compressor in communication-constrained non-IID scenarios, assigning different compression ratios to workers with different local dataset volumes.
+$\bar{\delta}$ of DAGC is equal to intial $\delta$ of DeCo-SGD.
+Accordion is the SOTA adaptive gradient compression algorithm, which sets small $\delta$ in the critical iterations and sets large $\delta$ if not.
+The large $\delta$ (as well as small $\delta$) is equal to the $\delta$ in DeCo-SGD under the constraint bandwidth (the half of the constraint bandwidth).
+DC2 adapts $\delta$ based on the network condition.
+We use DA2 in DC2 and $k_{0}$ is equal to intial $\delta$ of DeCo-SGD.
+The compressor we use in all baselines is Top-$k$, which transmits the largest $k$ elements of gradients.
+We select Top-$k$ for its empirical robustness and widespread adoption in DML.
+DGA first introduces delayed aggregation in FL to reduce the negative effect of the high latency.
+We set K=1K{=}1 in DGA to close the infrequent communication and $D$ equal to the initial $\tau$ of DeCo-SGD.
+CocktailSGD wang2023cocktailsgd uses the fixed $\tau{=}1$ and the compressor composed by Random-$k$ with $k{=}20\%$, Top-$k$ with $k{=}10\%$ and using $4$-bit quantization, as stated in the main body.
+Figure 6: The network bandwidth over time with an constraint of 100100Mbps.
+Figure 7: Training time comparison across different tasks.
+Network environment: We simulate challenging network environments (static high latency $\geq$ 100100ms, dynamic low bandwidth, constrained below 11Gbps).
+We measure the performance of our design under mulitple network conditions in Section $5.$ An example of dynamic bandwidth over time is shown in Fig.
+6.
 5.3 Comparison of Training Speed
 As illustrated in Fig.
-4, DeCo-SGD demonstrates significant improvements in training speed compared to D-SGD, Accordion, DGA, and CocktailSGD across all tasks.
-All experiments were conducted with $4$ worker nodes.
-Take ViT@ImageNet as an example, when the accuracy reaches 70%, DeCo-SGD achieves a 4.99×4.$9$ speedup444speed-up is defined as the ratio of the baseline method’s training time to that of DeCo-SGD.
-over D-SGD and a 1.34×1.34$\times$1.34 × speedup over CocktailSGD.
-At 80% accuracy, the speedup increases to 5.01×5.01$\times$5.01 × and 1.37×1.37$\times$1.37 ×, respectively.
+7, DeCo-SGD demonstrates significant improvements in training speed compared to D-SGD, Accordion, DGA, and CocktailSGD in all tasks.
+All experiments were conducted with $4$ worker nodes, under a fixed latency of 200200ms, with bandwidth fluctuating around 100100Mbps.
+Taking ViT@ImageNet as an example, when accuracy reaches 70%, DeCo-SGD achieves a $4.99\times$ speedup [^3] Speed-up is defined as the ratio of the baseline method’s training time to that of DeCo-SGD.
+over D-SGD and a $1.34\times$ speedup over CocktailSGD.
+At 80% accuracy, the speedup increases to $5.01\times$ and $1.37\times$, respectively.
 Compared with other baseline methods, the benefits of DeCo-SGD are particularly obvious, mainly for two reasons: (1) their parameter settings are suboptimal; (2) they cannot adapt to the dynamic network environment in real time.
 In contrast, DeCo-SGD utilizes theoretical guidance and runtime feedback to jointly optimize compression and latency, thereby achieving a more effective communication-computing trade-off throughout the training process.
-Figure 4: Training time comparison across different model-dataset pairs.
 5.4 Scalability Verification of DeCo-SGD
@@
 Fig.
-5 illustrates the performance of DeCo-SGD across different numbers of worker nodes, scaling from $n=4$ to $n=32$.
-All experiments were conducted under a fixed latency of $200$ms, with bandwidth fluctuating around $100$Mbps.
+8 illustrates the performance of DeCo-SGD across different numbers of worker nodes, scaling from n=4n{=}4 to n=32n{=}32.
 The results demonstrate that DeCo-SGD consistently maintains performance improvements as the number of nodes increases.
-Specifically, for GPT@Wikitext, DeCo-SGD achieves up to 3.77×3.$7$ and 1.21×1.21$\times$1.21 × speed-up over D-SGD and CocktailSGD, respectively, even at $n=32$.
-Similarly, for ViT@ImageNet, DeCo-SGD achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over CocktailSGD.
-These results highlight the adaptability of DeCo-SGD to large-scale and high-precision distributed training tasks.
+Specifically, for GPT@Wikitext, DeCo-SGD achieves up to $3.77\times$ and $1.21\times$ speed-up over D-SGD and CocktailSGD, respectively, even at n=32n{=}32.
+Similarly, for ViT@ImageNet, DeCo-SGD achieves up to $5.07\times$ speed-up over D-SGD and $1.37\times$ over CocktailSGD.
 Figure 8: Training time comparison of different methods when the worker node scales. DeCo-SGD demonstrates superior scalability and efficiency compared to other methods across both settings.
 5.5 DeCo-SGD under Different Bandwidths and Latency
-Table 1 presents the training times of different methods on GPT@Wikitext and ViT@ImageNet under varying network conditions, with all experiments conducted using $4$ worker nodes.
-The numbers in the table denote the training time (in seconds) required by each method to achieve a target perplexity ($22$ for the GPT model) or accuracy ($80$% for the ViT model).
-The results clearly demonstrate that under all network conditions, DeCo-SGD consistently outperforms other methods, achieving the lowest training times.
-This highlights its robustness against bandwidth and latency variations as well as its superior communication efficiency in challenging network environments.
-For instance, DeCo-SGD achieves up to 5.02×5.02$\times$5.02 × and 1.30×1.30$\times$1.30 × speed-ups over D-SGD and CocktailSGD, respectively.
-Table 1: Comparison of training time (s) under different methods and network conditions. $a$ is the average bandwidth and $b$ is the end-to-end latency. Detailed numbers are in the Appendix.
-Model @DataSet a(Gbps), b(s) D-SGD Accordion DGA CocktailSGD DeCo-SGD
-GPT @Wikitext $0.1,0.1$ $6396.95\pm 32.36$ $1955.28\pm 13.59$ $5506.78\pm 26.77$ $1525.21\pm 12.91$ $\mathbf{1306.29}\pm\mathbf{10.29}$
-$0.5,0.1$ $1571.37\pm 11.13$ $829.06\pm 6.65$ $1346.34\pm 10.75$ $712.93\pm 5.38$ $\mathbf{642.79}\pm\mathbf{5.74}$
-$0.1,1.0$ $7232.88\pm 35.62$ $2510.26\pm 22.73$ $5764.60\pm 29.69$ $1683.11\pm 14.66$ $\mathbf{1440.59}\pm\mathbf{11.87}$
-$0.5,1.0$ $2363.73\pm 13.87$ $1325.12\pm 10.90$ $1562.93\pm 11.89$ $1132.63\pm 9.91$ $\mathbf{997.25}\pm\mathbf{7.92}$
-ViT @ImageNet $0.1,0.1$ $1373.84\pm 10.58$ $440.14\pm 3.28$ $1183.68\pm 7.85$ $364.77\pm 1.02$ $\mathbf{283.06}\pm\mathbf{1.80}$
-$0.5,0.1$ $446.20\pm 3.39$ $274.60\pm 2.21$ $385.60\pm 1.57$ $219.39\pm 0.98$ $\mathbf{169.13}\pm\mathbf{0.47}$
-$0.1,1.0$ $1530.57\pm 9.89$ $604.67\pm 2.79$ $1287.03\pm 10.70$ $505.91\pm 2.56$ $\mathbf{397.16}\pm\mathbf{1.39}$
-$0.5,1.0$ $631.15\pm 3.33$ $373.79\pm 1.92$ $444.38\pm 2.18$ $278.57\pm 1.35$ $\mathbf{235.95}\pm\mathbf{0.88}$
+Table 3 presents the training times of different methods on GPT@Wikitext and ViT@ImageNet under varying network conditions, with all experiments conducted using $4$ worker nodes, using D-SGD, Accordion, DC2, DGA, CocktialSGD as baselines.
+Due to space limitations, we report the training time to reach a representative target metric for each task: $22$ for GPT and 8080% accuracy for ViT.
+Detailed convergence curves covering the entire training process are provided in the Appendix, which verify that DeCo-SGD consistently outperforms other methods, not just the specific points reported here.
+The advantage increases further as bandwidth decreases and network conditions worsen (bandwidth drops, latency increases).
+The results clearly demonstrate the robustness of our design against bandwidth and latency variations.
+The advantage of our design over baselines increases as network conditions worsen.
+We achieve up to $5.02\times$ and $1.30\times$ speed-ups over D-SGD and CocktailSGD, respectively.
+Notablely, in dynamic network environments, DC2 performs better than Accordion, but it does not address the problem of high latency.
+This shows that gradient compression alone cannot solve the problem.
+Table 3: Comparison of training time (s) under different methods and network conditions.
+$a$ is maximum bandwidth and $b$ is the end-to-end latency.
+The numbers in parentheses represent the speedup compared to DeCo-SGD.
+Model
+@DataSet
+a(Gbps),
+b(s)
+D-SGD Accordion DC2 DGA CocktailSGD DeCo-SGD
+GPT
+@Wikitext
+$0.1,0.1$ $6396.95\pm 32.36$ ($4.90\times$) $1955.28\pm 13.59$ ($1.50\times$) $1901.13\pm 11.45$ ($1.45\times$) $5506.78\pm 26.77$ ($4.22\times$) $1535.21\pm 12.91$ ($1.18\times$) $\mathbf{1306.29\pm 10.29}$
+$0.5,0.1$ $1571.37\pm 11.13$ ($2.44\times$) $829.06\pm 6.65$ ($1.29\times$) $801.72\pm 6.06$ ($1.24\times$) $1346.34\pm 10.75$ ($2.09\times$) $712.93\pm 5.38$ ($1.11\times$) $\mathbf{642.79\pm 5.74}$
+$0.1,1.0$ $7232.88\pm 35.62$ ($5.02\times$) $2510.26\pm 22.73$ ($1.74\times$) $2469.74\pm 22.98$ ($1.71\times$) $5764.60\pm 29.69$ ($4.00\times$) $1683.11\pm 14.66$ ($1.17\times$) $\mathbf{1440.59\pm 11.87}$
+$0.5,1.0$ $2363.73\pm 13.87$ ($2.37\times$) $1325.12\pm 10.90$ ($1.33\times$) $1321.25\pm 10.31$ ($1.32\times$) $1562.93\pm 11.89$ ($1.57\times$) $1132.63\pm 9.91$ ($1.14\times$) $\mathbf{997.25\pm 7.92}$
+ViT
+@ImageNet
+$0.1,0.1$ $1373.84\pm 10.58$ ($4.85\times$) $440.14\pm 3.28$ ($1.55\times$) $415.69\pm 3.14$ ($1.46\times$) $1183.68\pm 7.85$ ($4.18\times$) $364.77\pm 1.02$ ($1.29\times$) $\mathbf{283.06\pm 1.80}$
+$0.5,0.1$ $446.20\pm 3.39$ ($2.64\times$) $274.60\pm 2.21$ ($1.62\times$) $253.36\pm 2.17$ ($1.49\times$) $385.60\pm 1.57$ ($2.28\times$) $219.39\pm 0.98$ ($1.30\times$) $\mathbf{169.13\pm 0.47}$
+$0.1,1.0$ $1530.57\pm 9.89$ ($3.85\times$) $604.67\pm 2.79$ ($1.52\times$) $579.03\pm 2.53$ ($1.45\times$) $1287.03\pm 10.70$ ($3.24\times$) $505.91\pm 2.56$ ($1.27\times$) $\mathbf{397.16\pm 1.39}$
+$0.5,1.0$ $631.15\pm 3.33$ ($2.67\times$) $373.79\pm 1.92$ ($1.58\times$) $368.29\pm 1.83$ ($1.56\times$) $444.38\pm 2.18$ ($1.88\times$) $278.57\pm 1.35$ ($1.18\times$) $\mathbf{235.95\pm 0.88}$
+Table 4: Comparison of training time (s) under different methods and network conditions in non-IID scenarios.
+Model
+@DataSet
+a(Gbps),
+b(s)
+D-SGD DAGC DC2 DGA CocktailSGD DeCo-SGD
+GPT
+@Wikitext
+$0.1,0.1$ $7276.34\pm 49.33$ ($5.13\times$) $2283.53\pm 12.55$ ($1.61\times$) $2244.73\pm 12.21$ ($1.58\times$) $6108.13\pm 35.23$ ($4.31\times$) $1732.25\pm 12.58$ ($1.22\times$) $\mathbf{1417.55\pm 11.23}$
+$0.5,0.1$ $1785.64\pm 12.55$ ($2.41\times$) $969.11\pm 7.12$ ($1.30\times$) $917.54\pm 7.42$ ($1.23\times$) $1515.61\pm 12.48$ ($2.04\times$) $815.52\pm 6.45$ ($1.10\times$) $\mathbf{741.35\pm 5.73}$
+$0.1,1.0$ $8179.46\pm 54.32$ ($4.76\times$) $2841.85\pm 20.07$ ($1.65\times$) $2827.36\pm 20.91$ ($1.64\times$) $6517.52\pm 34.66$ ($3.79\times$) $2018.44\pm 14.65$ ($1.17\times$) $\mathbf{1718.41\pm 13.46}$
+$0.5,1.0$ $2736.48\pm 19.64$ ($2.31\times$) $1576.29\pm 12.98$ ($1.32\times$) $1570.48\pm 12.07$ ($1.32\times$) $1775.52\pm 13.29$ ($1.50\times$) $1459.16\pm 11.63$ ($1.23\times$) $\mathbf{1186.33\pm 9.27}$
+ViT
+@ImageNet
+$0.1,0.1$ $1273.93\pm 9.32$ ($4.64\times$) $567.73\pm 3.73$ ($2.06\times$) $537.85\pm 3.14$ ($1.95\times$) $1324.88\pm 10.26$ ($4.82\times$) $368.39\pm 1.82$ ($1.34\times$) $\mathbf{274.76\pm 1.45}$
+$0.5,0.1$ $434.16\pm 2.15$ ($2.34\times$) $364.22\pm 1.80$ ($1.93\times$) $353.70\pm 1.79$ ($1.91\times$) $467.97\pm 2.10$ ($2.52\times$) $224.09\pm 1.38$ ($1.21\times$) $\mathbf{185.90\pm 0.68}$
+$0.1,1.0$ $1724.74\pm 10.47$ ($4.22\times$) $672.44\pm 3.21$ ($1.64\times$) $644.02\pm 3.14$ ($1.57\times$) $1367.39\pm 8.25$ ($3.34\times$) $555.26\pm 3.11$ ($1.36\times$) $\mathbf{409.18\pm 2.48}$
+$0.5,1.0$ $677.23\pm 4.21$ ($2.78\times$) $489.41\pm 2.21$ ($2.00\times$) $478.93\pm 1.85$ ($1.96\times$) $599.93\pm 2.83$ ($2.46\times$) $283.89\pm 1.62$ ($1.17\times$) $\mathbf{243.52\pm 1.02}$
+5.6 DeCo-SGD under non-IID Scenarios
+Federated learning or other data silo scenarios encounter non-IID problems, leading to accuracy degradation niid2020icml.
+We compare DeCo-SGD with other algorithms in non-IID settings.
+To evaluate performance under these conditions, we simulate the label imbalance and quantity imbalance using a Dirichlet distribution with a concentration parameter of $0.5$ niidbench.
+Notably, for these experiments, we replace the Accordion with DAGC compared to Section 5.5, a more competitive baseline specifically designed for non-IID scenarios.
+In non-IID scenarios, we replace $\phi$ with $\phi^{\prime}=\phi/\delta$ according to Remark 1.
+Despite the increased training uncertainty and instability inherent in non-IID data, DeCo-SGD demonstrates exceptional adaptability.
+It achieves speed-ups ranging from $1.23\times$ to $2.06\times$ over the strong baseline DAGC, significantly outperforming the gains observed in IID settings.
+In non-IID scenarios, our design achieves up to $5.13\times$ and $1.36\times$ speed-ups over D-SGD and CocktailSGD.
 6 Related Work
-Gradient Compression Strategy.
-Gradient compression methods are generally categorized into quantization alistarh2017qsgd ; bernstein2018signsgd ; seide20141 , sparsity aji2017sparse ; lin2017deep ; stich2018sparsified , hybrid approaches dryden2016communication ; jiang2018sketchml , and low-rank strategies vogels2019powersgd .
-Traditional schemes often rely on fixed compression parameters, which limit their adaptability to varying network and training conditions.
-To address this, recent works have introduced dynamic and adaptive strategies.
-DC2 abdelmoniem2021dc2 adjusts communication based on real-time network dynamics, while DAGC lu2023dagc assigns compression aggressiveness based on each worker’s data volume.
-However, neither of them considers end-to-end latency.
-FedCG jiang2023heterogeneity selects representative clients and adapts compression ratios according to their heterogeneity and capacity changes.
-L-GreCo markov2024greco formulates the per-layer compression allocation as a constrained optimization problem solved via dynamic programming to balance model size and error.
-However, neither of the above two methods takes into account the dynamic changes of the network environment.
-Asynchronism.
-FedBuff nguyen2022federated uses a method of buffering asynchronous aggregation to improve communication efficiency and model performance.
-However, the buffering mechanism may increase memory overhead and delay the incorporation of fresh gradients.
-Mishchenko et al.mishchenko2022asynchronous rigorously proved theoretically and validated empirically that asynchronous SGD can achieve convergence comparable to mini-batch SGD, even in the presence of arbitrary gradient delays.
-CO2 sun2024co2 introduces local update and single-step asynchronous communication mechanisms to achieve complete overlap between communication and computing, but it assumes homogeneous device capabilities.
-To further address heterogeneity in node speed, Shadowheart SGD tyurinshadowheart dynamically adjusts the local update frequency and communication frequency of each node based on the different computing and communication speeds of each working node, thereby achieving the optimal time complexity in an asynchronous setting.
+Communication-efficient and adaptive training.
+To alleviate the communication bottlenecks in distributed training, gradient compression has become a standard practice.
+While early works focus on static compression strategies (e.g., fixed quantization bits or sparsity ratios), recent research has shifted towards dynamic and adaptive strategies enhance robustness against system heterogeneity or network dynamics.
+For instance, DAGC lu2023dagc assigns different compression ratios based on the local data volume of each worker to address communication problems in communication-constrained non-IID scenarios.
+FedCG Wu2021FedCGLC selects representative clients and adapts compression ratios according to their heterogeneity and capacity changes to make the system robust to device heterogeneity.
+HCEF HCEF theoretically analyzes the impact of local update frequency and gradient compression on the convergence bound in Cooperative Federated Edge Learning.
+It develops an online algorithm to dynamically determine these two parameters to minimize training latency and energy communication.
+HGC hu2024practical compresses both uplink and downlink communications in federated learning, removing redundancy in the training process to greatly reduce data transfer while keeping the model performance.
+DC2 dc2 adaptively controls the gradient compression ratio based on network delays, effectively managing network variations to strike a better balance between accelerating distributed training and preserving the model accuracy.
+These methods, however, are largely designed for bandwidth efficiency and remain insensitive to the fixed latency.
+This limitation leaves model performance vulnerable in challenging network environments where high latency coexists with limited bandwidth.
+Latency-tolerant training.
+To enhance the tolerance of training to high latency, relaxation of the strict synchronization barrier is a common approach.
+Fully asynchronous methods are primarily designed to address the straggler effect caused by heterogeneous device capabilities.
+For example, FedAC zhang2024fedac incorporates look-ahead momentum to correct updates in buffered aggregation, while FedFa xu2024fedfa utilizes an update buffer to enable fully asynchronous updates.
+To address computational heterogeneity, Shadowheart SGD tyurinshadowheart dynamically adjusts local update and communication frequencies, and P2P-AFL sad2025towards proposes a decentralized model fusion mechanism.
+While these approaches effectively reduce waiting times, However, the fully asynchronous strategy may introduce extremely high staleness, leading to model accuracy degradation.
+It is also unsuitable for WAN-based training, which has relatively fixed latency.
+To address this, DGA DGA introduces delayed aggregation into FL, attempting to solve the fixed end-to-end latency issue in FL.
+However, it does not account for communication bottlenecks.
+CO2 sun2024co2 introduces local update and delayed aggregation with $\tau{=}1$ to achieve complete overlap between communication and computing.
+But it focuses on high-speed interconnects and is not suitable for bad network conditions, which require adaptive adjustment of the staleness to overlap communication and computation.
+Distributed training in WANs.
+WAN-based training faces the challenges of high latency and low, varying bandwidth.
+Early efforts, such as MMOptree liu2022accelerating and NETSTORM li2024accelerating, optimize the logical topology to reduce hop counts.
+However, topology optimization alone cannot fully eliminate the communication bottlenecks and high end-to-end latency.
+StellaTrain lim2024accelerating integrates cache-aware compression and sparse optimization for geo-distributed clusters.
+However, StellaTrain lacks rigorous theoretical guidance regarding the trade-off between compression rate and staleness, relying instead on empirical tuning.
+Consequently, it may fail to maintain optimal convergence rates under the extreme and fluctuating network conditions typical of WANs.
 7 Conclusion
@@
 Building on these insights, we propose DeCo-SGD, an adaptive optimizer that dynamically adjusts compression and staleness based on network conditions.
-Our design achieves up to 5.07×5.07$\times$5.07 × speed-up over D-SGD and 1.37×1.37$\times$1.37 × over the SOTA method.
+Our design achieves up to $5.07\times$ speed-up over D-SGD and $1.37\times$ over the SOTA method.
 The code will be released upon publication.
 Appendix A Notation List
-Table 2: Notation list.
-Notation Description
-$n$ number of workers
-$\mathbf{T}$ total number of iterations
-$f_{i}(\cdot)$ the local loss function of worker $i$
-$f(\cdot)$ the global loss function, i.e., $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$
-$\delta$ compression ratio ($0<\delta\leq 1$)
-$\mathbf{C}_{\delta}(\cdot)$ the sparsification compressor with compression ratio $\delta$
-$\mathbf{x}^{i}_{t}$ the local model parameter in the $t$-th iteration of the $i$-th worker
-$\mathbf{x}_{t}$ the global model parameter in the $t$-th iteration
-$\mathbf{g}^{i}_{t}$ the stochastic gradient of $\textbf{x}_{t}$ of the worker $i$
-$\mathbf{g}_{t}$ the average of $\mathbf{g}^{i}_{t}$, i.e., $\mathbf{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}^{i}_{t}$
-$\gamma$ the stepsize
-$\tau$ the delayed staleness
-$\mathbf{e}_{t}^{i}$ the local error term of the $i$-th worker at the $t$-th iteration
-$L$ L-smoothness
-$\mu$ $\mu$-strongly convexity if $f_{i}$ is strongly convex
-$\zeta_{i}$ the data heterogenority of worker $i$
-$\zeta$ the global data heterogenority, i.e., $\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{2}$
-$\xi^{i}$ the stochastic gradient noise of worker $i$
-$\sigma^{2}$ the upper variance bound of the stochastic gradient noise
-$a$ the bandwidth (bits/s)
-$b$ the network end-to-end latency (s)
-$S_{g}$ the size of the gradient (bits)
-$T_{comp}$ computation time per iteration (s)
-$T_{avg}$ average end-to-end training time per iteration (s)
-$T_{sum}$ total end-to-end training time (s)
-$TS_{k}$ End time of the k-th computation(s)
-$TM_{k}$ End time of the k-th transmission(s)
-$TC_{k}$ End time of the k-th communication(s)
 Appendix B Detailed proof
@@
 If $c_{1},c_{2}\in\mathbb{R}^{d}$ then the Jensen’s inequality is: For all $\rho>0$, we have
-$\lVert c_{1}+c_{2}\rVert^{2}\leq(1+\rho)\lVert a\rVert^{2}+(1+\rho^{-1})\lVert b\rVert^{2}.$ (12)
+$\lVert c_{1}+c_{2}\rVert^{2}\leq(1+\rho)\lVert c_{1}\rVert^{2}+(1+\rho^{-1})\lVert c_{2}\rVert^{2}.$ (12)
 This can be written as:
-$\langle c_{1},c_{2}\rangle^{2}\leq\rho\lVert a\rVert^{2}+\rho^{-1}\lVert b\rVert^{2}.$ (13)
+$\langle c_{1},c_{2}\rangle^{2}\leq\rho\lVert c_{1}\rVert^{2}+\rho^{-1}\lVert c_{2}\rVert^{2}.$ (13)
 Lemma 2 (The nature of Top-$k$).
@@
 $\mathbb{E}_{\mathbf{C}_{\delta}}\lVert\mathbf{C}_{\delta}(\mathbf{x})-\mathbf{x}\rVert^{2}\leq(1-\delta)\lVert\mathbf{x}\rVert^{2}.$ (14)
-Lemma 3 (Lemma 27 of the work [19]).
+Lemma 3 (Lemma 27 of the work [stich2020communication]).
 Let $(r_{t})_{t\geq 0}$ and $(s_{t})_{t\geq 0}$ be sequences of positive numbers satisfying
@@
 steps are sufficient.
-Lemma 4 (Lemma 25 of the work [19]).
+Lemma 4 (Lemma 25 of the work [stich2020communication]).
 Let $(r_{t})_{t\geq 0}$ and $(s_{t})_{t\geq 0}$ be sequences of positive numbers satisfying
-$r_{t+1}\leq(1-\min\{\gamma A,F\})r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
-for some positive constants $A,B>0$, $C,D\geq 0$, and for constant step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$, and for parameter $0<F\leq 1$.
-Then there exists a constant step-size $\gamma\leq\frac{1}{E}$ such that
-$\frac{B}{W_{T}}\sum_{t=0}^{T}w_{t}s_{t}+\min\left\{A,\frac{F}{\gamma}\right\}r_{T+1}$ $\leq r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}(T+1)\right]$
-$\quad+\frac{2C\ln\tau}{A(T+1)}+\frac{D\ln^{2}\tau}{A^{2}(T+1)^{2}}$
-for $w_{t}:=(1-\min\{\gamma A,F\})^{-(t+1)}$, $W_{T}:=\sum_{t=0}^{T}w_{t}$ and
-$\tau=\max\left\{\exp[1],\min\left\{\frac{A^{2}r_{0}(T+1)^{2}}{C},\frac{A^{3}r_{0}(T+1)^{3}}{D}\right\}\right\}$
+$r_{t+1}\leq(1-\gamma A)r_{t}-B\gamma s_{t}+C\gamma^{2}+D\gamma^{3},$
+for some positive constants $A,B>0$, $C,D\geq 0$, and for constant step-sizes $0<\gamma\leq\frac{1}{E}$, for $E\geq 0$. Then we have
+$\frac{B}{W_{T}}\sum_{t=0}^{T}w_{t}s_{t}+Ar_{T+1}\leq r_{0}E\exp\left[-\frac{A}{E}(T+1)\right]+\frac{2C\ln\tau}{A(T+1)}+\frac{D\ln^{2}\tau}{A^{2}(T+1)^{2}},$
+for $w_{t}:=(1-\gamma A)^{-(t+1)}$, $W_{T}:=\sum_{t=0}^{T}w_{t}$ and $\tau=\max\left\{\exp[1],\min\left\{\frac{A^{2}r_{0}(T+1)^{2}}{C},\frac{A^{3}r_{0}(T+1)^{3}}{D}\right\}\right\}$.
 Remark 6.
 Lemma  4 establishes a bound of the order
-$\tilde{\mathcal{O}}\left(r_{0}\left(E+\frac{A}{F}\right)\exp\left[-\min\left\{\frac{A}{E},F\right\}T\right]+\frac{C}{AT}+\frac{D}{A^{2}T^{2}}\right),$
+$\tilde{\mathcal{O}}\left(r_{0}E\exp\left[-\frac{A}{E}\cdot T\right]+\frac{C}{AT}+\frac{D}{A^{2}T^{2}}\right),$
 that decreases with $T$. To ensure that this expression is less than $\epsilon$,
-$T=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}\log\frac{1}{\epsilon}+\frac{E}{A}\log\frac{1}{\epsilon}\right)=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{1}{F}+\frac{E}{A}\right)$
+$T=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{E}{A}\log\frac{1}{\epsilon}\right)=\tilde{\mathcal{O}}\left(\frac{C}{A\epsilon}+\frac{\sqrt{D}}{A\sqrt{\epsilon}}+\frac{E}{A}\right)$
 steps are sufficient.
@@
 If $\gamma\leq\frac{1}{4L}$, then we have
-$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+\gamma L^{2}\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+2\gamma L^{2}\mathbb{E}\left\|B_{t}\right\|^{2}+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
+$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+\gamma L^{2}(2\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+\mathbb{E}\left\|B_{t}\right\|^{2})+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
 Proof. We begin with the definition of $\hat{\mathbf{x}}_{t+1}$ and have
 $\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)$ $=\mathbb{E}f\left(\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right)$ (16)
-$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\frac{L}{2}\mathbb{E}\left\|\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right\|^{2}$
-$\leq\mathbb{E}f\left(\tilde{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
+$\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\frac{L}{2}\mathbb{E}\left\|\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right\|^{2}$
+$\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
 The first line is due to $\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}$.
@@
 1 and the third line is due to Eq.
-3.
+2.
 We estimate $A_{1}$ as
@@
 $=\gamma(\mathbb{E}\left\langle\nabla f\left(\tilde{\mathbf{x}}_{t}\right)-\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle+\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right)-\nabla f\left(\tilde{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle-\mathbb{E}\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{g}_{t}\rangle)$
-$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{c_{2}\gamma L^{2}}{2}\mathbb{E}\lVert B_{t}\rVert^{2}+\frac{\gamma}{2c_{2}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}.$
-The last line is due to Jensen’s inequality (i.e., Eq. 13). We let $c_{1}=2$, $c_{2}=4$. By bringing Eq. LABEL:A-1 into Eq. 16, we complete the proof.
+$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{c_{2}\gamma L^{2}}{2}\mathbb{E}\lVert B_{t}\rVert^{2}+\frac{\gamma}{2c_{2}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}$
+$\leq-\frac{5\gamma}{8}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\gamma L^{2}(2\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
+The third line is due to Jensen’s inequality (i.e., Eq. 13). In the fourth line, we let $c_{1}=4$, $c_{2}=2$. Bringing Eq. LABEL:A-1 into Eq. 16, we complete the proof.
 Lemma 6.
-Define $\{\tilde{B}_{t}\}_{t\geq 0}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{4L\sqrt{\tau}}$, we have
+Define $\{\tilde{B}_{t}\}_{t\geq 0}$ as in Eq. 6. Then with $\gamma\leq\frac{1}{4L\tau}$, we have
 $\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{16L^{2}}\sum_{t=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{\tau\gamma^{2}(1+T)\sigma^{2}}{n}.$
-Proof. This lemma follows from the Lemma 20 in the previous work [14].
+Proof.
+$\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $=\gamma^{2}\mathbb{E}\lVert\sum_{j=1}^{\tau}(\nabla f(\mathbf{x}_{t-j})+\frac{1}{n}\sum_{i=1}^{n}\xi^{i})\rVert^{2}$
+$\leq\gamma^{2}(\tau\sum_{j=1}^{\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-j})\rVert^{2}+\frac{1}{n^{2}}\cdot n\tau\sigma^{2}).$
+The inequality is due to Jensen’s inequality and Eq. 2. Then we have
+$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $\leq\gamma^{2}\tau\sum_{t=0}^{T}\sum_{j=1}^{\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-j})\rVert^{2}+\frac{\gamma^{2}\tau\sigma^{2}(1+T)}{n}$ (18)
+$\leq\gamma^{2}\tau\cdot\tau\sum_{t=0}^{T-\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{\gamma^{2}\tau\sigma^{2}(1+T)}{n}$
+$\leq\gamma^{2}\tau^{2}\sum_{t=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{\gamma^{2}\tau\sigma^{2}(1+T)}{n}.$
+The second line is due to the fact that the term $\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}$ appears at most $\tau$ times. Let $\gamma\leq\frac{1}{4L\tau}$, we complete the proof.
 Lemma 7.
 Define $\{B_{t}\}_{t\geq 0}$ as in Eq.
-3.2.
-Then by using the relative compressor, with $\gamma\leq\frac{1}{4L\sqrt{\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, we have
-$\sum_{t=0}^{T}\mathbb{E}\lVert B_{t}\rVert^{2}\leq\frac{1}{4L^{2}}\mathbb{E}\sum_{t=0}^{T}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi(T+1)(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
+6.
+Then by using the relative compressor, with $\gamma\leq\frac{1}{4LZ\sqrt{\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, we have
+$\sum_{t=0}^{T}\mathbb{E}\lVert B_{t}\rVert^{2}\leq\frac{1}{4L^{2}}\sum_{t=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi(T+1)(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
 Proof.
-$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\frac{(1-\delta)(1+\rho)}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}$
-$+\gamma^{2}(1-\delta)((1+\rho^{-1})\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+(1+\rho^{-1})\zeta^{2}+\sigma^{2}).$
-This inequality is due to Lemma 2, Eq. 4, 3,12.
-By unrolling the recurrence and let $\rho=\frac{2-\delta}{\delta}$, we can derive the upper bound:
-$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
-$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
-$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
-$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{T}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
-Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we have
+$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\frac{(1-\delta)(1+\rho)}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}$ (19)
+$\quad+\gamma^{2}(1-\delta)((1+\rho^{-1})Z^{2}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+(1+\rho^{-1})\zeta^{2}+\sigma^{2})$
+$\leq\frac{1}{n}(1-\frac{\delta}{2})\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau}^{i}\rVert^{2}+\gamma^{2}(1-\delta)(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-\tau})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
+The first inequality is due to Lemma 2, Eq.
+3, 2 and 12.
+In the second line, we let $\rho=\frac{\delta}{2-\delta}$, having $1+\rho^{-1}=\frac{2}{\delta}$ and $(1-\delta)(1+\rho)=1-\frac{\delta}{2-\delta}<1-\frac{\delta}{2}$.
+By unrolling the recurrence, we can derive the upper bound:
+$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
+$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
+$\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}\sum_{i=0}^{T}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
+$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{T}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
+Choosing $\frac{4\gamma^{2}(1-\delta)Z^{2}}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we have
 $\sum_{t=0}^{T}[\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}]\leq\frac{1}{4L^{2}}\sum_{i=0}^{T}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2\gamma^{2}(1-\delta)(1+T)}{\delta(1-\frac{\delta}{2})^{\tau}}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}).$
@@
 Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth.
-There exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\sqrt{\tau}},\frac{1}{4L\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
-$\begin{split}\mathcal{O}(&\frac{\sigma^{2}L}{n\epsilon^{2}}+\frac{L\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{L}{\epsilon}+\frac{L\sqrt{\tau}}{\epsilon}+\frac{\sqrt{\phi}L}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
+There exists a stepsize $\gamma\leq\min\{\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$, such that at most
+$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon}+\frac{Z\sqrt{\phi}}{\sqrt{\delta}\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
 iterations of DD-EF-SGD, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total iterations, chosen at random uniformly.
 Proof. Let $G_{t}:=\sum_{i=0}^{t}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{i}\right)\right\|^{2}$. According to above lemmas (Lemma 5, 6, 7), we have
-$\frac{\gamma}{2}G_{T}\leq\mathbb{E}f(\hat{\mathbf{x}}_{0})-\mathbb{E}f(\hat{\mathbf{x}}_{T+1})+\gamma L^{2}(\sum_{i=0}^{T}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}\lVert B_{t}\rVert^{2})+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}$
-$\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma}{4}G_{T}+\frac{\gamma}{8}G_{T}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+2\gamma^{3}L^{2}(T+1)(2\phi\frac{2}{\delta}\zeta^{2}+2\phi\sigma^{2}+\frac{2\tau\sigma^{2}}{n}).$
-Divide both sides of the inequality by $4\gamma(1+T)$, we have:
-$\frac{1}{8(1+T)}G_{T}\leq\frac{\mathbb{E}f(\mathbf{x}_{0})-f^{*}}{\gamma(T+1)}+\gamma\frac{L\sigma^{2}}{2n}+4\gamma^{2}L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n}).$
-According to Lemma 3, we let $B=1/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n})$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$.
-To ensure the right hand side of the above inequality is less than $\epsilon>0$, $T=\mathcal{O}(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon})r_{0}$ iterations are sufficient, which completes the proof.
+$\frac{\gamma}{2}G_{T}$ $\leq\mathbb{E}f(\hat{\mathbf{x}}_{0})-\mathbb{E}f(\hat{\mathbf{x}}_{T+1})+\gamma L^{2}(2\sum_{i=0}^{T}\lVert\tilde{B}_{t}\rVert^{2}+\sum_{i=0}^{T}\lVert B_{t}\rVert^{2})+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}$
+$\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma}{4}G_{T}+\frac{\gamma}{8}G_{T}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+\gamma^{3}L^{2}(T+1)(2\phi\frac{2}{\delta}\zeta^{2}+2\phi\sigma^{2}+\frac{2\tau\sigma^{2}}{n}).$
+Divide both sides of the inequality by $4\gamma(1+T)$, and we have:
+$\frac{1}{8(1+T)}G_{T}\leq\frac{\mathbb{E}f(\mathbf{x}_{0})-f^{*}}{\gamma(T+1)}+\gamma\frac{L\sigma^{2}}{2n}+2\gamma^{2}L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n}).$
+We do not believe that DD-EF-SGDwill degrade into D-SGD, meaning that $\tau=0$ and $\delta=1$ cannot coexist at the same time.
+In this way, we have $\gamma\leq\min\{\frac{1}{4L},\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}\leq\min\{\frac{1}{4L\tau},\frac{1}{4LZ\sqrt{\phi/\delta}}\}$.
+According to Lemma 3, we let $B=1/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=L^{2}(\frac{2\phi}{\delta}\zeta^{2}+\phi\sigma^{2}+\frac{\tau\sigma^{2}}{n})$, $E=\max\{4L\tau,4LZ\sqrt{\phi/\delta}\}$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$.
+To ensure the right hand side of the above inequality is less than $\epsilon>0$, $T=\mathcal{O}(\frac{C}{\epsilon^{2}}+\frac{\sqrt{D}}{\epsilon^{3/2}}+\frac{E}{\epsilon})\cdot r_{0}$ iterations are sufficient, which completes the proof.
 B.3 Proof of Theorem 2
@@
 Proof.
-$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-x^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
+$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-\mathbf{x}^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
 We estimate the term $A_{2}$ as
@@
 $\leq-\frac{\gamma\mu}{2}X_{t}-\gamma F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
-The second line is due to the reason that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{4L},\mu\leq L$, we have
-$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}+(2L\gamma^{2}-\gamma)F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})$
-$\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}f_{t}+6L\gamma(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2}).$
+The second line is due to the fact that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{4L},\mu\leq L$, we have
+$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}+(2L\gamma^{2}-\gamma)F_{t}+\gamma(4L+2\mu)(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})+\gamma^{2}\sigma^{2}$
+$\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+6L\gamma(\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\mathbb{E}\lVert B_{t}\rVert^{2})+\gamma^{2}\sigma^{2}.$
 This completes the proof.
 Lemma 9.
-Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $\tilde{B}_{t}$ as in Eq. 3.2. Then with $\gamma\leq\frac{1}{8\sqrt{L\tau}}$, we have
-$\sum_{t=0}^{T}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n},$
-where $W_{t}=\sum_{i=1}^{n}w_{i}$.
-Proof. This lemma follows from the Lemma 20 in the previous work [14].
+Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $\tilde{B}_{t}$ as in Eq. 6. Then with $\gamma\leq\frac{1}{4\sqrt{2}\ L\tau}$, we have
+$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}F_{t}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n},$
+where $W_{t}=\sum_{i=0}^{t}w_{i}$.
+Proof. Based on Eq. 18, when $\gamma\leq\frac{1}{4\sqrt{2}L\tau}$, we have
+$\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $\leq\frac{1}{128L^{2}\tau}\sum_{j=1}^{\tau}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t-j})\rVert^{2}+\frac{\tau\gamma^{2}\sigma^{2}}{n}$
+$\leq\frac{1}{64L\tau}\sum_{j=1}^{\tau}F_{t-j}+\frac{\tau\gamma^{2}\sigma^{2}}{n}.$
+The second inequality is due to the assumption of L-smoothness $\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert\leq 2LF_{t}$.
+We observe that when $\tau\geq 1$, we have $w_{t}=(1-\frac{\mu\gamma}{2})^{-(t+1)}$, $1-\frac{\mu\gamma}{2}\geq 1-\frac{\mu}{8\sqrt{2}\ L\tau}\geq 1-\frac{1}{8\sqrt{2}\ \tau}\leq(1+\frac{1}{2\tau})^{-1}$, thus having $w_{t}=w_{t-j}(1-\frac{\mu\gamma}{2})^{-j}\leq w_{t-j}(1+\frac{1}{2\tau})^{j}\leq w_{t-j}(1+\frac{1}{2\tau})^{\tau}\leq w_{t-j}\exp[1/2]\leq 2w_{t-j}$.
+Based on this inequality, we have
+$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}\frac{w_{t}}{2\tau}\sum_{j=1}^{\tau}F_{t-j}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n}$
+$\leq\frac{1}{32L}\sum_{t=0}^{T}\frac{1}{\tau}\sum_{j=1}^{\tau}w_{t-j}F_{t-j}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n}$
+$\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}F_{t}+\frac{W_{T}\tau\gamma^{2}\sigma^{2}}{n}.$
 Lemma 10.
 Let $w_{t}=(1-\frac{\gamma\mu}{2})^{-(t+1)}$ and define $B_{t}$ as in Eq.
-3.2.
-Then by using the relative compressor, with $\gamma\leq\frac{1}{16\sqrt{L\phi}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$ we have
-$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma^{2}\phi W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
-where $W_{t}=\sum_{i=1}^{n}w_{i}$.
-proof. Denote $E_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$, and we have:
-$w_{t}E_{t}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
-$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
-$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}w_{i}(1+\frac{\delta}{4})^{t-i}(\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2})$
-$+\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}$
-$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{i=0}^{t}(1-\frac{\delta}{4})^{t-i}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}]$
-$\sum_{t=0}^{T}w_{t}E_{t}$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{j=0}^{\infty}(1-\frac{\delta}{4})^{j}\sum_{i=0}^{T}w_{i}\frac{2}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
-$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{4}{\delta}\sum_{i=0}^{T}w_{i}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})].$
-Choosing $\frac{4\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{4L^{2}}$, we complete the proof.
+6.
+Then by using the relative compressor, with $\gamma\leq\frac{1}{16LZ\sqrt{2\phi/\delta}}$, where $\phi=\frac{1-\delta}{\delta(1-\frac{\delta}{2})^{\tau}}$ we have
+$\sum_{t=0}^{T}w_{t}\mathbb{E}\lVert B_{t}\rVert^{2}$ $\leq\frac{1}{32L}\sum_{t=0}^{T}w_{t}F_{t}+2\gamma^{2}\phi W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2}),$
+where $W_{t}=\sum_{i=0}^{t}w_{i}$.
+proof. We denote $E_{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\lVert e_{t-\tau+1}^{i}\rVert^{2}$, and based on Eq. 19, then we have:
+$w_{t}E_{t}$ $\leq\gamma^{2}\frac{(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t-\tau}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
+$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(1-\frac{\delta}{2})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\frac{2}{\delta}\zeta^{2}+\sigma^{2})$
+$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}(1-\frac{\delta}{2})^{t-i}w_{i}(1+\frac{\delta}{4})^{t-i}(\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2})$
+$\quad+\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}$
+$\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{i=0}^{t}(1-\frac{\delta}{4})^{t-i}w_{i}\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{i=0}^{t}w_{t}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})(1-\frac{\delta}{2})^{t-i}],$
+$\sum_{t=0}^{T}w_{t}E_{t}$ $\leq\frac{\gamma^{2}(1-\delta)}{(1-\frac{\delta}{2})^{\tau}}[\sum_{j=0}^{\infty}(1-\frac{\delta}{4})^{j}\sum_{i=0}^{T}w_{i}\frac{2Z^{2}}{\delta}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+\sum_{j=0}^{\infty}(1-\frac{\delta}{2})^{j}W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
+$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{4Z^{2}}{\delta}\sum_{i=0}^{T}w_{i}\mathbb{E}\lVert\nabla f(\mathbf{x}_{i})\rVert^{2}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})]$
+$\leq\frac{2\gamma^{2}(1-\delta)}{\delta(1-\frac{\delta}{2})^{\tau}}[\frac{8LZ^{2}}{\delta}\sum_{i=0}^{T}w_{i}F_{i}+W_{T}(\frac{2}{\delta}\zeta^{2}+\sigma^{2})].$
+Choosing $\frac{16LZ^{2}\gamma^{2}(1-\delta)}{\delta^{2}(1-\frac{\delta}{2})^{\tau}}\leq\frac{1}{32L}$, we complete the proof.
 Theorem 2 (Convex convergence rate of DD-EF-SGD).
 Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-convex.
-Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4L},\frac{1}{8L\sqrt{\tau}},\frac{1}{16L\sqrt{\phi}}\}$, such that at most
-$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L}{\mu}+\frac{\sqrt{L\tau}}{\mu}+\frac{\sqrt{L\phi}}{\mu})\end{split}$
-iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\min\left\{\frac{\mu\gamma}{2},\frac{\delta}{4}\right\})^{-t}$.
+Then there exists a stepsize $\gamma\leq\min\{\frac{1}{4\sqrt{2}\ L\tau},\frac{1}{16LZ\sqrt{2\phi/\delta}}\}$, such that at most
+$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L(\frac{\phi\zeta^{2}}{\delta}+(\phi+\frac{\tau}{n})\sigma^{2})}}{\mu\epsilon^{1/2}}+\frac{L(\tau+Z\sqrt{\phi/\delta})}{\mu})\end{split}$
+iterations of DD-EF-SGD, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\frac{\mu\gamma}{2})^{-t}$.
 Proof. By substituting Lemma 5, 6 into Lemma 4, we can get
@@
 According to Lemma 4, we complete the proof.
-B.4 Proof of Theorem 3
+B.4 Proof of Convergence rate of D-SGD with delayed aggregation
+The update rule of D-SGD with delayed aggregation is: $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t-\tau}^{i}$. Based on the virtual sequence, we have
+$\tilde{B}_{t}=\frac{\gamma}{n}\sum_{i=1}^{n}\sum_{j=1}^{\tau}\mathbf{g}_{t-\tau}^{i}{,}\quad\hat{\mathbf{x}}_{t+1}=\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\mathbf{g}_{t}^{i}.$ (20)
+Simlar to Section B.2, we have following lemmas.
+Lemma 11.
+Let $\left\{\mathbf{x}_{t},\hat{\mathbf{x}}_{t},\tilde{B}_{t}\right\}_{t\geq 0}$ be defined in D-SGD with delayed aggregation (Eq.
+20), with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth.
+If $\gamma\leq\frac{3}{2L}$, then we have
+$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\frac{\gamma}{2}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{t}\right)\right\|^{2}+2\gamma L^{2}\mathbb{E}\left\|\tilde{B}_{t}\right\|^{2}+\frac{\gamma^{2}L\sigma^{2}}{2n}.$
+Proof. We begin with the definition of $\hat{\mathbf{x}}_{t+1}$ and have
+$\mathbb{E}f\left(\hat{\mathbf{x}}_{t+1}\right)$ $=\mathbb{E}f\left(\hat{\mathbf{x}}_{t}-\frac{\gamma}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\right)$ (21)
+$\leq\mathbb{E}f\left(\hat{\mathbf{x}}_{t}\right)-\underbrace{\gamma\mathbb{E}\left\langle\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle}_{:=A_{1}}+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{L\gamma^{2}\sigma^{2}}{2n}.$
+We estimate $A_{1}$ as
+$A_{1}$ $=\gamma(\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right)-\nabla f\left(\hat{\mathbf{x}}_{t}\right),\mathbf{g}_{t}\right\rangle-\mathbb{E}\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{g}_{t}\rangle)$ (22)
+$\leq\frac{c_{1}\gamma L^{2}}{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\frac{\gamma}{2c_{1}}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}-\gamma\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}$
+$\leq-\frac{7\gamma}{8}\mathbb{E}\lVert\nabla f(\mathbf{x}_{t})\rVert^{2}+2\gamma L^{2}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}.$
+The second line is due to Jensen’s inequality (i.e., Eq. 13). We let $c_{1}=4$ in the third line. We bring $A_{1}$ into Eq. 21 and complete the proof.
+Theorem 4 (Non-convex convergence rate of D-SGD with delayed aggregation).
+Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth. Then there exists a stepsize $\gamma\leq\frac{1}{4L\tau}$, such that at most
+$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\frac{\tau}{n}\sigma^{2}}}{\epsilon^{3/2}}+\frac{\tau}{\epsilon})\cdot L(f(\mathbf{x}_{0})-f^{*})\end{split}$
+iterations of D-SGD with delayed aggregation, it holds $\mathbb{E}\lVert\nabla f(\mathbf{x}_{out})\rVert^{2}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{\mathbf{T}-1}\right\}$, where $\mathbf{T}$ denotes the total iterations, chosen at random uniformly.
+Proof. Let $G_{t}:=\sum_{i=0}^{t}\mathbb{E}\left\|\nabla f\left(\mathrm{x}_{i}\right)\right\|^{2}$. According to above lemmas (Lemma 11, 6), we have
+$\frac{3\gamma}{8}G_{T}\leq\mathbb{E}f(\mathbf{x}_{0})-f^{*}+\frac{\gamma^{2}L(T+1)\sigma^{2}}{2n}+\gamma^{3}L^{2}(T+1)\frac{2\tau\sigma^{2}}{n}.$
+In this way, we have $\gamma\leq\min\{\frac{3}{2L},\frac{1}{4L\tau}\}\leq\frac{1}{4L\tau}$.
+According to Lemma 3, we let $B=3/8$, $C=\frac{L\sigma^{2}}{2n}$, $D=\frac{L^{2}\tau\sigma^{2}}{n}$, $E=4L\tau$, $r_{0}=\mathbb{E}f(\mathbf{x}_{0})-f^{*}$, completing the proof.
+Lemma 12.
+Let $\left\{\mathbf{x}_{t},\hat{\mathbf{x}}_{t},\tilde{B}_{t}\right\}_{t\geq 0}$ be defined in D-SGD with delayed aggregation (Eq.
+20), with gradient oracle $\left\{\mathbf{g}_{t}\right\}_{t\geq 0}$ and let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-strongly convex.
+If $\gamma\leq\frac{1}{2L}$, then we have
+$X_{t+1}$ $\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+\gamma^{2}\sigma^{2}+5\gamma L\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}.$
+Proof.
+$X_{t+1}$ $=X_{t}+\gamma^{2}\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{t}^{i}\rVert^{2}-\underbrace{2\gamma\mathbb{E}\left\langle\hat{\mathbf{x}}_{t}-\mathbf{x}^{*},\mathbf{g}_{t}\right\rangle}_{:=A_{2}}.$
+We estimate the term $A_{2}$ as
+$A_{2}$ $=-2\gamma(\mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\mathbf{x}^{*}\right\rangle-\ \mathbb{E}\left\langle\nabla f\left(\mathbf{x}_{t}\right),\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\right\rangle)$
+$\leq-\gamma\mu\mathbb{E}\lVert\mathbf{x}_{t}-\mathbf{x}^{*}\rVert^{2}-2\gamma F_{t}+\frac{1}{2}\gamma F_{t}+4\gamma L\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$
+$\leq-\frac{\gamma\mu}{2}X_{t}-\frac{3}{2}\gamma F_{t}+\gamma(4L+\mu)\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}.$
+The second line is due to the fact that $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth and $\mu$-strongly convex. Then choosing $\gamma\leq\frac{1}{2L},\mu\leq L$, we have
+$X_{t+1}\leq(1-\frac{\gamma\mu}{2})X_{t}-\frac{\gamma}{2}F_{t}+5L\gamma\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}+\gamma^{2}\sigma^{2}.$
+This completes the proof.
+Theorem 5 (Convex convergence rate of D-SGD with delayed aggregation).
+Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $L$-smooth and $\mu$-strongly convex. Then there exists a stepsize $\gamma\leq\frac{1}{4\sqrt{2}\ L\tau}$, such that at most
+$\begin{split}\mathcal{O}(&\frac{\sigma^{2}}{n\mu\epsilon}+\frac{\sqrt{L\frac{\tau}{n}\sigma^{2}}}{\mu\epsilon^{1/2}}+\frac{L\tau}{\mu})\end{split}$
+iterations of D-SGD with delayed aggregation, it holds $\mathbb{E}f(\mathbf{x}_{out})-f^{*}\leq\epsilon$, and $\mathbf{x}_{out}=\mathbf{x}_{t}$ denotes an iterate $\mathbf{x}_{t}\in\left\{\mathbf{x}_{0},\ldots,\mathbf{x}_{T-1}\right\}$, selected probabilistically based on $(1-\frac{\mu\gamma}{2})^{-t}$.
+Proof. By substituting Lemma 9 into Lemma 12, we can get
+$X_{T}\leq w_{T}X_{0}-\frac{\gamma}{2}\sum_{i=0}^{T}w_{t}F_{t}+\frac{W_{T}\gamma^{2}\sigma^{2}}{n}+5\gamma L\sum_{i=0}^{T}w_{t}\mathbb{E}\lVert\tilde{B}_{t}\rVert^{2}$
+$\leq w_{T}X_{0}-\frac{11\gamma}{32}\sum_{i=0}^{T}w_{t}F_{t}+W_{T}\gamma^{2}\sigma^{2}+5LW_{T}\gamma^{3}\frac{\tau\sigma^{2}}{n}.$
+According to Lemma 4, we complete the proof.
+B.5 Proof of Theorem 3
 Theorem 3 (Estimation of $T_{avg}$ in DD-EF-SGD).
-Assume that throughout $t$ iterations where $\tau\ll t\ll\mathbf{T}$, the constants $a$, $b$, $\delta$, and $\tau$ remain fixed.
-Then the average iteration time $T_{avg}:=\frac{TC_{t}}{t}$ satisfies the following approximation:
-$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1},\frac{\delta S_{g}}{a},T_{\text{comp}}\right\}},$ (18)
-where the approximation error is upper bounded by $O(\frac{1}{t})$.
+Under Assumptions from 5 to 8, the average iteration time $T_{avg}$ satisfies the following approximation:
+$T_{avg}\approx\max{\left\{\frac{T_{\text{comp}}+b+\delta\cdot S_{g}/a}{\tau+1},\frac{\delta\cdot S_{g}}{a},T_{\text{comp}}\right\}},$ (23)
+where the approximation error is upper bounded by $O(\frac{1}{E})$, where $E$ is the analysis window.
 proof. Based on the assumption, the iterative formula of $TS_{k}$ derived from the relationship between computation and communication can be written as :
-$\left\{\begin{matrix}TC_{k}=TM_{k}+b\\ TS_{k+1}=T_{\text{comp}}+max\left\{TC_{k-\tau},TS_{k}\right\}\\ TM_{k+1}=\frac{\delta S_{g}}{a}+max\left\{TM_{k},TS_{k+1}\right\}\end{matrix}\right.\quad,k=0,1,2...\quad.$ (19)
+$\left\{\begin{matrix}TC_{k}=TM_{k}+b\\ TS_{k+1}=T_{\text{comp}}+max\left\{TC_{k-\tau},TS_{k}\right\}\\ TM_{k+1}=\frac{\delta S_{g}}{a}+max\left\{TM_{k},TS_{k+1}\right\}\end{matrix}\right.\quad,k=0,1,2.$$\quad.$ (24)
 with initial values $TS_{0}=TM_{0}=0$ and $TC_{k}=0$ for all $k\leq 0$.
@@
 It follows that:
-$TC_{t}=TM_{t}+b=\frac{\delta S_{g}}{a}+tT_{\text{comp}}+b,$
+$TC_{E}=TM_{E}+b=\frac{\delta S_{g}}{a}+ET_{\text{comp}}+b,$
 and thus the approximation error satisfies:
-$\left|TC_{t}-tT_{\text{comp}}\right|=\frac{\delta S_{g}}{a}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
+$\left|TC_{E}-ET_{\text{comp}}\right|=\frac{\delta S_{g}}{a}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
 Case 2: Communication-Dominated Regime
@@
 with:
-$\left|TC_{t}-t\cdot\frac{\delta S_{g}}{a}\right|=T_{\text{comp}}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
+$\left|TC_{E}-E\cdot\frac{\delta S_{g}}{a}\right|=T_{\text{comp}}+b\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
 Finally, we verify the existence of $N_{1}>\tau$ such that $TC_{N_{1}-\tau}\geq TS_{N_{1}}$.
@@
 This periodic property can be verified by explicitly analyzing the first $(2\tau+2)$ terms of the sequence $\{(TS_{k},TM_{k})\}$. We have:
-{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\text{for }i=1,\dots,\tau+1;$ (20)
-{$TS_{i}&=\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\\ TM_{i}&=2\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\end{aligned}\right.\quad\text{for }i=\tau+2,\dots,2\tau+2.$
+$\left\{\begin{aligned} TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=\frac{\delta S_{g}}{a}+iT_{\text{comp}},\end{aligned}\right.\quad\text{for }i=1,\dots,\tau+1;$ (25)
+$\left\{\begin{aligned} TS_{i}&=\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\\ TM_{i}&=2\frac{\delta S_{g}}{a}+b+(i-\tau)T_{\text{comp}},\end{aligned}\right.\quad\text{for }i=\tau+2,\dots,2\tau+2.$
 Let us define the index group:
-$A_{k}$ :=$:=\left\{(TM_{k(\tau+1)+d},TS_{k(\tau+1)+d})\mid d=1,\dots,\tau+1\right\},$
-$A_{k}+a$ :=$:=\left\{(TM_{k(\tau+1)+d}+a,TS_{k(\tau+1)+d}+a)\mid d=1,\dots,\tau+1\right\},\quad\forall a\in\mathbb{R}.$
+$A_{k}$ $=\left\{(TM_{k(\tau+1)+d},TS_{k(\tau+1)+d})\mid d=1,\dots,\tau+1\right\},$
+$A_{k}+a$ $=\left\{(TM_{k(\tau+1)+d}+a,TS_{k(\tau+1)+d}+a)\mid d=1,\dots,\tau+1\right\},\quad\forall a\in\mathbb{R}.$
 By rewriting the iterative update rule in the compact form $A_{k}=F(A_{k-1})$, it follows from translation invariance that:
@@
 $A_{k+1}=A_{k}+\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right),\quad\forall k\in\mathbb{N}^{*}.$
-Let $t=m(\tau+1)+d$ for $d=1,\dots,\tau+1$, then the cumulative time $TC_{t}$ satisfies:
-$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+dT_{\text{comp}}+\frac{\delta S_{g}}{a}+b,$
-$\left|TC_{t}-t\left(\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)\right|$ $=d\left(T_{\text{comp}}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+\frac{\delta S_{g}}{a}+b$
+Let $E=m(\tau+1)+d$ for $d=1,\dots,\tau+1$, then the cumulative time $TC_{E}$ satisfies:
+$TC_{E}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+dT_{\text{comp}}+\frac{\delta S_{g}}{a}+b,$
+$\left|TC_{E}-E\left(\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)\right|$ $=d\left(T_{\text{comp}}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+\frac{\delta S_{g}}{a}+b$
 $\leq\tau T_{\text{comp}}\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
-This confirms that $TC_{t}$ deviates from its average rate by at most a bounded constant, further supporting the regularity and predictability of the update pattern in this intermediate-delay regime.
+This confirms that $TC_{E}$ deviates from its average rate by at most a bounded constant, further supporting the regularity and predictability of the update pattern in this intermediate-delay regime.
 Case 4: Periodic Structure under Computation-Dominated Regime
@@
 Analogously to Case 3, we examine the first $(3\tau+3)$ elements of the sequence $\left\{(TS_{k},TM_{k})\right\}$ to validate the periodic structure:
-{$TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=i\frac{\delta S_{g}}{a}+T_{\text{comp}},\end{aligned}\right.\quad$ $i=1,\dots,\tau+1;$
-{$TS_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-2)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-1)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=\tau+2,\dots,2\tau+2;$
-{$TS_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-3)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-2)\tfrac{\delta S_{g}}{a},\end{aligned}\right.\quad$ $i=2\tau+3,\dots,3\tau+3.$
+$\left\{\begin{aligned} TS_{i}&=iT_{\text{comp}},\\ TM_{i}&=i\frac{\delta S_{g}}{a}+T_{\text{comp}},\end{aligned}\right$$\quad$ $i=1,\dots,\tau+1;$
+$\left\{\begin{aligned} TS_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-2)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-\tau-1)\tfrac{\delta S_{g}}{a},\end{aligned}\right$$\quad$ $i=\tau+2,\dots,2\tau+2;$
+$\left\{\begin{aligned} TS_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-3)\tfrac{\delta S_{g}}{a},\\ TM_{i}&=2(b+T_{\text{comp}}+\tfrac{\delta S_{g}}{a})+T_{\text{comp}}+(i-2\tau-2)\tfrac{\delta S_{g}}{a},\end{aligned}\right$$\quad$ $i=2\tau+3,\dots,3\tau+3.$
 We define the grouping $A_{k}$ as in Case 3. From the recurrence observed above, we find that:
@@
 Thus, the $(\tau+1)$-periodicity is established, completing the proof for Case 4.
-Now, let $t=m(\tau+1)+d$ with $d\in\{1,\dots,\tau+1\}$. The cumulative computation time $TC_{t}$ satisfies:
-$TC_{t}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+d\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}+b,$
-$\left|TC_{t}-t\cdot\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right|$ $=d\left(\frac{\delta S_{g}}{a}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+T_{\text{comp}}+b$
+Now, let $E=m(\tau+1)+d$ with $d\in\{1,\dots,\tau+1\}$. The cumulative computation time $TC_{E}$ satisfies:
+$TC_{E}$ $=m\left(T_{\text{comp}}+b+\frac{\delta S_{g}}{a}\right)+d\cdot\frac{\delta S_{g}}{a}+T_{\text{comp}}+b,$
+$\left|TC_{E}-E\cdot\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right|$ $=d\left(\frac{\delta S_{g}}{a}-\frac{T_{\text{comp}}+b+\frac{\delta S_{g}}{a}}{\tau+1}\right)+T_{\text{comp}}+b$
 $\leq\tau\cdot\frac{\delta S_{g}}{a}\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}.$
@@
 By integrating the results from all four cases, we obtain a unified expression for the asymptotic average iteration time:
-$T^{\prime}_{\text{avg}}$ =$=\begin{cases}\dfrac{T_{\text{comp}}+\frac{\delta S_{g}}{a}+b}{\tau+1},&\tau<\tau_{\text{thres}},\\ \max\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},&\tau\geq\tau_{\text{thres}},\end{cases}\quad\text{with}\quad\left|TC_{t}-tT^{\prime}_{\text{avg}}\right|\leq b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\},$
+[…]
 where $T^{\prime}_{\text{avg}}:=\lim\limits_{k\to\infty}\frac{TS_{k}}{k}$ approximates the true average iteration time $T_{\text{avg}}$, and
@@
 with the absolute error bounded by:
-$\left|T_{\text{avg}}-T^{\prime}_{\text{avg}}\right|\leq\frac{b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}}{t}=O\left(\frac{1}{t}\right).$
+$\left|T_{\text{avg}}-T^{\prime}_{\text{avg}}\right|\leq\frac{b+\min\left\{T_{\text{comp}},\frac{\delta S_{g}}{a}\right\}}{E}=O\left(\frac{1}{E}\right).$
 Appendix C Addendum to Evaluation Experiments
-C.1 Experimental Environment
-The experiments are conducted on an Ubuntu $24.04$ LTS server with an Intel Xeon Gold $6230$ CPU and $3$ Nvidia A40404040 GPUs, each with $48$GB memory.
-The Python version is $3.10.16$, and other used libraries are all based on the Python version.
-We use PyTorch $2.5.1$ with CUDA $12.4$ as the ML toolkit.
-We use gloo as the communication backend.
-C.2 Detailed Experimental Settings
-Experiment tasks: The experimental setup encompasses three task types spanning computer vision (CV) and natural language processing (NLP), utilizing both traditional (CNN [29]) and state-of-the-art transformer-based models (ViT [30] and GPT [31]).
-The CNN architecture includes two convolutional layers followed by two fully connected layers, used widely in previous works [9, 20].
-The ViT base model features approximately 86 million parameters, while the smallest GPT variant contains 124 million parameters.
-To demonstrate the superiority of our design, we evaluate four representative tasks: FashionMNIST [32] and CIFAR-$10$ with CNN (denoted as CNN@FMNIST and CNN@CIFAR-$10$), ImageNet [33] using ViT (denoted as ViT@ImageNet) and Wikitext [34] on GPT (denoted as GPT@Wikitext).
-For CNN@FMNIST and CNN@CIFAR-$10$, the learning rate is $0.01$, and the batch size is $32$.
-For ViT@ImageNet and GPT@Wikitext, the learning rates are $0.01$ and $0.1$, respectively, with batch sizes of $40$ and $5$.
-Baselines: We compare D-SGD, Accordion, DGA, and vanilla DD-EF-SGD as baselines.
-D-SGD is the case $\tau=0,\delta=1$, and While it does not suffer from accuracy degradation, it converges the slowest in high-latency and low-bandwidth scenarios when measured in terms of end-to-end time.
-Accordion [35] is the SOTA adaptive gradient compression algorithm, which sets small $\delta$ in the critical iterations and sets large $\delta$ if not [9].
-The compressor we use in all baselines is Top-$k$, which transmits the largest $k$ elements of gradients.
-The reason why we select Top-$k$ is its strong empirical performance, widespread adoption in DML, and straightforward implementation.
-DGA is a state-of-the-art algorithm originally designed for Federated Learning [18], which first introduces delay staleness to reduce the negative effect of the high latency on Federated Learning.
-We set $K=1$ in DGA to close the infrequent communication.
-CocktailSGD [21] uses the fixed $\tau$ and $\delta$.
-The hyperparameters are chosen by DeCo-SGD with $E=\infty$.
-Network environment: We simulated challenging network environments (static high latency $\geq$ $100$ms, dynamic low bandwidth, average bandwidth $\leq$ $1$Gbps), where the dynamic volatility of the network can be referred to the following figure, which gives the network environment in the given environment (Section 5.2), and we discuss the superiority of the DeCo-SGD algorithm in different network conditions (Section 5.4).
-C.3 Network Conditions of Section 5.2
+C.1 Compression ratio and stelaness curves in Section 5.3
 As shown in Fig.
-LABEL:fig:network, the network conditions exhibit dynamic bandwidth fluctuations across all tasks (CNN@FMNIST, CNN@CIFAR-10, ViT@ImageNet, and GPT@Wikitext), and the compression ratio $\delta$ is adaptively adjusted over time under a fixed 200 ms transmission delay with $\tau=2$.
-C.4 Detailed Curves of Section 5.3 to Verify the Scalability of DeCo-SGD
-The detail of Fig. 5 in the main text of the paper corresponds to Fig. 9 in the Appendix.
+9, the network conditions exhibit dynamic bandwidth fluctuations across all tasks (CNN@FMNIST, CNN@CIFAR-10, ViT@ImageNet, and GPT@Wikitext), and the compression ratio $\delta$ is adaptively adjusted over time under a fixed 200 ms transmission delay with $\tau=2$.
+(a) ResNet@CIFAR-10
+(b) ViT@ImageNet
+(c) GPT@Wikitext
+Figure 9: $\delta$ varying over time for tasks: ResNet@CIFAR-10, ViT@ImageNet, and GPT@Wikitext, with $\tau=2$.
+C.2 Detailed Curves of Section 5.4 to Verify the Scalability of DeCo-SGD
+Fig. 10 provide the detailed accuracy/ppl curves in Fig. 8,
 (a) ViT@ImageNet $n=4$
@@
 Bottom row: Perplexity over time for GPT@Wikitext.
-C.5 Detailed Results of Section 5.4
-Fig. 10 in the Appendix provides a visual counterpart to the data presented in Table 1 in the main text. Table 3 extends more details in Table 1
-Table 3: Comparison of training time (s) under different methods and network conditions.
-$a$ is the average bandwidth and $b$ is the end-to-end latency.
-$\tau^{*}$ and $\delta^{*}$ are caculated by DeCo-SGD.
-The numbers in parentheses represent the speedup compared to DeCo-SGD.
-Model @DataSet a(Gbps), b(s) $\tau^{*}$, $\delta^{*}$ D-SGD Accordion DGA CocktailSGD DeCo-SGD
-GPT @Wikitext $0.1,0.1$ $2,0.02$ $6396.95$ (4.90×4.90$\times$4.90 ×) $1955.28$ (1.50×1.50$\times$1.50 ×) $5506.78$ (4.22×4.$2$) $1525.21$ (1.17×1.17$\times$1.17 ×) $\mathbf{1306.29}$
-$0.5,0.1$ $2,0.10$ $1571.37$ (2.44×2.$4$) $829.06$ (1.29×1.29$\times$1.29 ×) $1346.34$ (2.09×2.09$\times$2.09 ×) $712.93$ ($1\times1$) $\mathbf{642.79}$
-$0.1,1.0$ $3,0.02$ $7232.88$ (5.02×5.02$\times$5.02 ×) $2510.26$ (1.74×1.74$\times$1.74 ×) $5764.60$ (4.00×4.$0$) $1683.11$ (1.17×1.17$\times$1.17 ×) $\mathbf{1440.59}$
-$0.5,1.0$ $3,0.10$ $2363.73$ (2.37×2.37$\times$2.37 ×) $1325.12$ (1.33×1.$3$) $1562.93$ (1.57×1.57$\times$1.57 ×) $1132.63$ (1.14×1.14$\times$1.14 ×) $\mathbf{997.25}$
-ViT @ImageNet $0.1,0.1$ $2,0.03$ $1373.84$ (4.85×4.85$\times$4.85 ×) $440.14$ (1.55×1.$5$) $1183.68$ (4.18×4.18$\times$4.18 ×) $364.77$ (1.29×1.29$\times$1.29 ×) $\mathbf{283.06}$
-$0.5,0.1$ $2,0.14$ $446.20$ (2.64×2.64$\times$2.64 ×) $274.60$ (1.62×1.62$\times$1.62 ×) $385.60$ (2.28×2.28$\times$2.28 ×) $219.39$ (1.30×1.30$\times$1.30 ×) $\mathbf{169.13}$
-$0.1,1.0$ $3,0.03$ $1530.57$ (3.85×3.85$\times$3.85 ×) $604.67$ (1.52×1.52$\times$1.52 ×) $1287.03$ (3.24×3.24$\times$3.24 ×) $505.91$ (1.27×1.27$\times$1.27 ×) $\mathbf{397.16}$
-$0.5,1.0$ $3,0.14$ $631.15$ (2.67×2.67$\times$2.67 ×) $373.79$ (1.58×1.58$\times$1.58 ×) $444.38$ (1.88×1.$8$) $278.57$ (1.18×1.18$\times$1.18 ×) $\mathbf{235.95}$
-(a) ViT@ImageNet bandwidth=$100$Mbps latency=$100$ms
-(b) ViT@ImageNet bandwidth=$100$Mbps latency=$1000$ms
-(c) ViT@ImageNet bandwidth=$500$Mbps latency=$100$ms
-(d) ViT@ImageNet bandwidth=$500$Mbps latency=$1000$ms
-(e) GPT@Wikitext bandwidth=$100$Mbps latency=$100$ms
-(f) GPT@Wikitext bandwidth=$100$Mbps latency=$1000$ms
-(g) GPT@Wikitext bandwidth=$500$Mbps latency=$100$ms
-(h) GPT@Wikitext bandwidth=$500$Mbps latency=$1000$ms
+C.3 Detailed Results of Section 5.5
+Fig. 11 provides a visual counterpart to the data presented in Table 3 in the main text.
+(a) ViT@ImageNet bandwidth=100100Mbps latency=100100ms
+(b) ViT@ImageNet bandwidth=100100Mbps latency=10001000ms
+(c) ViT@ImageNet bandwidth=500500Mbps latency=100100ms
+(d) ViT@ImageNet bandwidth=500500Mbps latency=10001000ms
+(e) GPT@Wikitext bandwidth=100100Mbps latency=100100ms
+(f) GPT@Wikitext bandwidth=100100Mbps latency=10001000ms
+(g) GPT@Wikitext bandwidth=500500Mbps latency=100100ms
+(h) GPT@Wikitext bandwidth=500500Mbps latency=10001000ms
 Figure 11: Convergence performance across different numbers of worker nodes ($n=4,8,16,32$) for five distributed optimization methods.
@@
 Bottom row: Perplexity over time for GPT@Wikitext.
-Appendix D Others
-D.1 UsedResources and Licenses
-All external assets used in this paper are properly credited, and their licenses and terms of use have been explicitly acknowledged and respected.
-Specifically, we make use of several open-source resources as follows:
-  1. 1.
-    ViT model, released under the MIT License.
-  2. 2.
-    GPT2 model from Hugging Face, released under the MIT License.
-  3. 3.
-    docker-tc, used to simulate network conditions, released under the MIT License.
-All of these assets were used strictly for research purposes in accordance with their respective licenses.
-We have cited the corresponding authors or repositories in our references where appropriate and ensured full compliance with their terms of use.
-D.2 Limitations and Boarder Impact
-Limitations: (1) Although DeCo-SGD currently supports only the D-SGD and its variants, we believe that D-SGD is applicable to a wide range of tasks, including NLP, CV and Objective Detection.
-If D-SGD is properly parameterized, its performance is no worse than popular optimizers in LLM training such as Adam [52].
-(2) Device heterogeneity (different bandwidth/latency per node) is not considered, as current theoretical frameworks (based on convex optimization) cannot analyze such scenarios without prohibitive parameter search costs.
-Nevertheless, our approach already achieves $80\%$ throughput gains over SOTA optimizer [21] in dynamic networks, leaving limited room for further improvement via heterogeneity optimization.
-Boarder Impact: Our work establishes theoretical foundations for LLM pre-training in low-bandwidth, high-latency networks.
-While employing lossy training techniques, we provide theoretical guarantees of training efficiency.
-This could significantly reduce the infrastructure costs for LLM pre-training, promoting broader participation in AI research and development.
+C.4 Experiment Results in non-IID settings
+The non-IID partition strategy is distribution-based label imbalance, which allocates a proportion of samples of each label according to Dirichlet distribution, with the concentration parameter set to 0.5.
+The results demonstrate that under non-IID settings, DeCo-SGD consistently outperforms other methods.
+C.5 Experiment Results with other sparsifiers
+We compare DeCo-SGD with other optimizers under another sparsifier Random-K, which randomly selects k% gradient elements to transmit.
+The results below show that DeCo-SGD consistently outperforms other methods with other sparsifiers.
+Table 5: Comparison of training time for DeCo-SGD and other optimizers under the Random-K sparsifier, which randomly selects k% gradient elements to transmit.
+Model
+@DataSet
+a(Gbps),
+b(s)
+D-SGD DAGC Accordion DC2 DGA CocktailSGD DeCo-SGD
+GPT
+@Wikitext
+$0.1,0.1$ $7276.34\pm 49.33$ ($5.13\times$) $2283.53\pm 12.55$ ($1.61\times$) $2246.34\pm 12.24$ ($1.59\times$) $2244.73\pm 12.21$ ($1.58\times$) $6108.13\pm 35.23$ ($4.31\times$) $1732.25\pm 12.58$ ($1.22\times$) $\mathbf{1417.55\pm 11.23}$
+$0.5,0.1$ $1785.64\pm 12.55$ ($2.41\times$) $969.11\pm 7.12$ ($1.30\times$) $921.27\pm 7.56$ ($1.24\times$) $917.54\pm 7.42$ ($1.23\times$) $1515.61\pm 12.48$ ($2.04\times$) $815.52\pm 6.45$ ($1.10\times$) $\mathbf{741.35\pm 5.73}$
+$0.1,1.0$ $8179.46\pm 54.32$ ($4.76\times$) $2841.85\pm 20.07$ ($1.65\times$) $2812.31\pm 20.35$ ($1.63\times$) $2827.36\pm 20.91$ ($1.64\times$) $6517.52\pm 34.66$ ($3.79\times$) $2018.44\pm 14.65$ ($1.17\times$) $\mathbf{1718.41\pm 13.46}$
+$0.5,1.0$ $2736.48\pm 19.64$ ($2.31\times$) $1576.29\pm 12.98$ ($1.32\times$) $1566.14\pm 12.11$ ($1.32\times$) $1570.48\pm 12.07$ ($1.32\times$) $1775.52\pm 13.29$ ($1.50\times$) $1459.16\pm 11.63$ ($1.23\times$) $\mathbf{1186.33\pm 9.27}$
+ViT
+@ImageNet
+$0.1,0.1$ $1273.93\pm 9.32$ ($4.64\times$) $567.73\pm 3.73$ ($2.06\times$) $545.18\pm 3.45$ ($1.98\times$) $537.85\pm 3.14$ ($1.95\times$) $1324.88\pm 10.26$ ($4.82\times$) $368.39\pm 1.82$ ($1.34\times$) $\mathbf{274.76\pm 1.45}$
+$0.5,0.1$ $434.16\pm 2.15$ ($2.34\times$) $364.22\pm 1.80$ ($1.93\times$) $347.22\pm 1.73$ ($1.87\times$) $353.70\pm 1.79$ ($1.91\times$) $467.97\pm 2.10$ ($2.52\times$) $224.09\pm 1.38$ ($1.21\times$) $\mathbf{185.90\pm 0.68}$
+$0.1,1.0$ $1724.74\pm 10.47$ ($4.22\times$) $672.44\pm 3.21$ ($1.64\times$) $654.28\pm 3.09$ ($1.60\times$) $644.02\pm 3.14$ ($1.57\times$) $1367.39\pm 8.25$ ($3.34\times$) $555.26\pm 3.11$ ($1.36\times$) $\mathbf{409.18\pm 2.48}$
+$0.5,1.0$ $677.23\pm 4.21$ ($2.78\times$) $489.41\pm 2.21$ ($2.00\times$) $467.39\pm 2.07$ ($1.92\times$) $478.93\pm 1.85$ ($1.96\times$) $599.93\pm 2.83$ ($2.46\times$) $283.89\pm 1.62$ ($1.17\times$) $\mathbf{243.52\pm 1.02}$